### v1

# Benchmarking Cross-Domain Audio-Visual Deception Detection

Abstract
Automated deception detection is crucial for assisting humans in accurately assessing truthfulness and identifying deceptive behavior. Conventional contact-based techniques, like polygraph devices, rely on physiological signals to determine the authenticity of an individual’s statements. Nevertheless, recent developments in automated deception detection have demonstrated that multimodal features derived from both audio and video modalities may outperform human observers on publicly available datasets. Despite these positive findings, the generalizability of existing audio-visual deception detection approaches across different scenarios remains largely unexplored. To close this gap, we present the first cross-domain audio-visual deception detection benchmark, that enables us to assess how well these methods generalize for use in real-world scenarios. We used widely adopted audio and visual features and different architectures for benchmarking, comparing single-to-single and multi-to-single domain generalization performance. To further exploit the impacts using data from multiple source domains for training, we investigate three types of domain sampling strategies, including domain-simultaneous, domain-alternating, and domain-by-domain for multi-to-single domain generalization evaluation. Furthermore, we proposed the Attention-Mixer fusion method to improve performance, and we believe that this new cross-domain benchmark will facilitate future research in audio-visual deception detection. Protocols and source code are available at https://github.com/Redaimao/cross_domain_DD.
Index Terms:
audio-visual, multimodal deception detection, cross-domain, generalization.
1 Introduction
Figure 1: Typical samples from different publicly deception detection datasets: Real Life Trials [1], Bag of Lies [2], MU3D [3], and Box of Lies [4]. The samples in each row are from different datasets while those in each column are with different modalities (visual vs. audio) and ground truth labels (i.e., truthful vs. deceptive). It can be seen that serious domain shifts (e.g., resolution/illumination/pose in visual faces and pitch/loudness/noise in audio) occur among these datasets.
Audio-visual deception detection involves utilizing AI techniques and algorithms to automatically detect deceptive behavior in speech and facial movements [5, 6, 7, 8]. Deception detection has a significant impact on various real-world applications such as law enforcement [9], healthcare [10], and business [11]. It has the potential to prevent fraud, improve security measures, and enhance trust and confidence. A reliable deception detection tool can support more accurate decision-makings.
Traditional deception detection is often a contact-based method. It assesses whether someone is telling the truth or not by monitoring physiological responses like skin conductance and heart rate [12]. Experts’ Behavioral observation and analysis are another technique that evaluates changes in a person’s body language, speech patterns, and eye movements [13, 14]. However, such an assessment can be time-consuming and require significant expertise to perform accurately.
Recently, the development of automated deception detection systems using AI and machine learning techniques has gained significant attention as the existing methods have limitations in terms of reliability, accuracy, and scalability. Various multimodal datasets have been introduced, including real-life trials from court scenes [1], lab-based setups [2, 3], and game show scenarios [4]. These datasets provide a wide variety of deceptive samples from different domains, enabling researchers to examine the effectiveness of AI models on deception detection. Based on these datasets, progress has been made in deception detection techniques within specific domains [7, 6, 15]. Recent studies have utilized rich visual and audio features [16, 17, 18], such as Mel Spectrogram, emotional states, and facial action units, to enhance the performance of deception detection tasks.
However, there remains a substantial research gap that needs to be addressed. Specifically, fewer studies have explored the cross-domain issue, despite the presence of significant domain shifts in public deception detection datasets. As shown in Figure 1, domain shifts are observed in both audio and visual modalities from publicly available datasets. The generalizability of the models is critical for practical applications. Therefore, such domain shifts need to be investigated in order to develop deception detection models that can be generalized across different contexts. Additionally, effective methods must be proposed to alleviate the domain shift issue by fusing both audio and visual features in a meaningful way. Addressing these issues can benefit automated deception detection systems in improving generalizability in real-world applications.
To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets. Our benchmarking approach utilizes widely adopted audio and visual features, and we compare the single-to-single domain performance and multi-to-single domain generalization using different architectures. Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing. To further enhance performance, we propose an Attention-Mixer fusion method based on MLP-Mixer [19]. This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings. Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues. In summary, our main contributions include:
  • Introducing a new benchmark for evaluating the generalization capacity of AI models using audio and visual features across different domains.
  • Comparing the single-to-single domain and multi-to-single domain generalization using different architectures.
  • Providing three domain sampling strategies i.e., domain simultaneous, domain alternating, and domain-by-domain, to conduct multi-to-single cross-domain testing.
  • Proposing the Attention-Mixer fusion method to enhance performance.
In the rest of the paper, Sec. 2 provides a review of related psychological studies on cues to deception and multimodal deception detection works. Sec. 3 introduces our benchmarking approach and fusion method. Sec. 4 provides the cross-domain benchmark results and fusion results. Finally, conclusions and future works are given in Sec. 5.
2 Related work
2.1 Cues to Deception
The research on using behavioral cues for deception has gradually become active over the past few decades. Psychological researchers have published a large number of works on the analysis of cues to deception [20, 21, 22]. Among the studied behavioral cues, verbal and nonverbal cues were preferred as humans may behave differently between lying and telling the truth. DePaulo et al. [23] studied and reported experimental results on 158 cues to deception. They revealed that, in general, people who tell lies are less forthcoming and less convincing than those who tell the truth. Liars usually talk about fewer details and make fewer spontaneous corrections. They also sound less involved but more vocally tense. Through the study, the researchers statistically found that liars often press their lips, repeat words, raise their chins, and show less genuine smiles. The results show that some behavioral cues do potentially appear in deception and are even more pronounced when liars are more motivated to cheat.
Levine et al. [24] reviewed the status quo and provided a new perspective on the theories of deception. They pointed out that lying usually happens when problematic information is involved. It is critical to understand the verbal content in the context. Vrij et al. [25] realized that interviewers play a vital role in eliciting and enhancing cues to deceit. The authors proposed the “interviewing to detect deception” technique to open a new path in the deception detection research field. They argued that different psychological states can be exploited by adopting appropriate interview techniques of liars and truth-tellers. Hirschberg et al. [26] proposed a method to distinguish deceptive from non-deceptive speeches by a large corpus. They also conducted experiments using acoustic, lexical, and speaker-dependent features, which showed improved performance by combining multiple feature sets. Warren et al. [27] conducted experiments to investigate the relationship between affective facial expressions to deception. The results indicated that leaked emotions with the incongruous intended message can provide useful cues to deception, which supported the nonverbal leakage theory [28, 29].
Figure 2: Main network and method. (a) Model architecture. Visual modality includes face and behavior inputs. Audio modality includes Mel Spectrogram input. The features obtained by the respective encoders. The fusion methods include score fusion and feature fusion. (b) Domain sampling strategies. Domain-simultaneous: each batch consists of samples from multiple sources. Domain-alternating: each batch is alternatively sampled from multiple sources. Domain-by-domain: the batches are sampled from one source and then from another.
2.2 Multimodal Deception Detection
Recent works for deception detection usually use verbal and non-verbal features and propose effective fusion methods [30, 15, 16, 17]. For example, some works utilized facial features from RGB images to perform deception detection [15, 17, 5]. To capture facial movements, facial action units (AUs) were utilized. Other features, such as facial expression, were also adopted [6, 30]. Besides visual features, many works incorporated audio features to boost performance [7, 6, 15]. For example, Wu et al. [6] used MFCC (Mel-frequency Cepstral Coefficients) features and Karimi et al. [15] used raw audio. Most of the recent works mentioned above have considered multimodal fusion approaches that extract visual, audio, and text information to boost performance. In addtion to visual, audio and text, Karnati et al. [16] exploited physiological signals, i.e., EEG representations for deception detection.
To better fuse the multimodal features, different fusion methods were proposed, which can be broadly categorized into feature-level fusion and decision-level fusion. Specifically, feature-level fusion focused on producing better multimodal embeddings and used the linear layers to extract crossmodal dynamics [15, 17, 5, 6, 7]. Rather than that, decision-level fusion aimed to fuse multimodal dynamics at a late stage, to reduce computational complexity and learn good marginal representations [16, 7].
However, previous works on multimodal deception detection did not consider cross-domain issues that occur from one domain to another, which is the focus of this work.
3 Methodology
The mainstream architecture for audio-visual deception detection usually includes encoders for unimodal feature extraction and/or a fusion module. We follow the widely adopted architecture to build the benchmark on cross-domain audio-visual deception detection in this work. As shown in Fig. 2, audio and visual features are extracted from audio and visual encoders. The fusion module is performed based on audio and visual features. The fused feature is input to the classifier for classification. We build the benchmark for cross-domain generalization performance based on such network architecture with different encoders. We conducted single-to-single and multi-to-single evaluations where three domain sampling strategies included, i.e., domain-simultaneous, domain-alternating, and domain-by-domain.
3.1 Audio and Visual Feature Learning
To establish a benchmark for cross-domain audio-visual deception detection, we utilize widely adopted audio and visual features along with their respective encoders. Our approach treats audio and visual features as equally important, extracting different types of features simultaneously. As depicted in Fig. 1, this network structure offers several advantages: (1) flexibility in network selection: different audio or visual encoders can be effortlessly incorporated and compared in a fair manner, (2) adaptability: the addition or removal of specific modules and/or losses is straightforward. For instance, a fusion module can be inserted before classifiers, and (3) easy performance benchmarking: the system facilitates evaluating performance in various settings, such as score-level fusion and feature-level fusion. In this work, we focus on audio and visual modalities for deception detection. In particular, two kinds of visual features are extracted, i.e., face features from RGB face images and behavior features consist of AUs, affects, etc.
As shown in Fig. 1, given a detected RGB face image as input $X_{f}$, the deep features $F_{f}$ could be extracted via face encoder networks $\mathcal{E}_{f}$ (e.g., ResNet18 [31]). Similarly, behavior inputs such as the AU and/or affect features $X_{b}$ are encoded by OpenFace [32] or affect model (e.g., EmotionNet [33]) $\mathcal{E}_{b}$ to output behavior features $F_{b}$. Note that we regard both face frames and behavior features as the visual modality but differentiate them in this work as they have different types of information and representations. Given audio input $X_{a}$ (either Mel Spectrogram [34] or waveforms), audio features $F_{a}$ are extracted through audio encoder $\mathcal{E}_{a}$. The corresponding classification heads for face frames ($\mathcal{H}_{f}$), behavior features ($\mathcal{H}_{b}$), and audio features ($\mathcal{H}_{a}$) output the prediction logits $\hat{Y}_{f}$, $\hat{Y}_{b}$, and $\hat{Y}_{a}$, respectively. The fusion head $\mathcal{G}$ takes $F_{f}$, $F_{b}$, and $F_{a}$ as input. $\mathcal{G}$ is determined by the actual fusion method, e.g., liner layer, transformer layers, MLP, etc. The output logit of $\mathcal{G}$ is denoted by $\hat{Y}_{g}$. Therefore, the audio and visual learning process can be denoted as follows:
$\small\begin{split}F_{f}&=\mathcal{E}_{f}(X_{f}),\hat{Y}_{f}=\mathcal{H}_{f}(F_{f}),\\ F_{b}&=\mathcal{E}_{b}(X_{b}),\hat{Y}_{b}=\mathcal{H}_{b}(F_{b}),\\ F_{a}&=\mathcal{E}_{a}(X_{a}),\hat{Y}_{a}=\mathcal{H}_{a}(F_{a}),\\ \hat{Y}_{g}&=\mathcal{G}(F_{f},F_{b},F_{a}).\end{split}$ (1)
Loss Function. For deception detection ground truth $Y$, where $Y=0$ for truthful and $Y=1$ for deception, the binary cross-entropy loss (BCE) is adopted. The loss for each sample with a certain modality or fused prediction can be denoted as
$\mathcal{L}_{m}=-(Ylog(\hat{Y}_{m})+(1-Y)log(1-\hat{Y}_{m})),$ (2)
where $m\in\{f,b,a,g\}$, $\hat{Y}_{m}$ is the corresponding prediction logits. In other words, the BCE loss is calculated separately for each type of modality and/or its fused feature depending on whether a sample has any face frames, visual inputs, or audio inputs. The overall loss function can be described as follows:
$\mathcal{L}=\frac{1}{N}\sum_{i=i}^{N}\left(\sum_{m={f,b,a}}\mathcal{L}_{m,i}+\lambda\mathcal{L}_{g,i}\right),$ (3)
where $N$ is the number of data samples and $\lambda$ is a trade-off parameter between modality loss and fusion loss. $\lambda$ is set to 0.5 in our experiments.
3.2 Cross-domain Generalization
We benchmark the cross-domain generalization on the deception detection task. First, we introduce the notations and definitions in this section. A domain is composed of data that are sampled from a distribution (dataset), which can be denoted as $\mathcal{S}=\{(X;Y)_{i}\}_{i=1}^{N}\sim P_{S}$, where $X=(X_{f},X_{b},X_{a})$, $X_{f},X_{b},X_{a}$ represent samples of face frames, behavior, and audio modalities, respectively. $Y$ denotes the label, and $P_{S}$ denotes the joint distribution of the input samples and the output label. In this paper, for simplicity, we follow similar definitions in [35, 36] to treat each dataset as an individual domain due to their obvious distribution gaps, but more fine-grained intra-domain factors would be explored in future work. For domain generalization, $M$ source domains (training datasets) are given, i.e., $\mathcal{S}_{train}=\{S_{j}|j=1,\cdots,M\}$, where $\mathcal{S}_{j}=\{(X;Y)_{i}\}_{i=1}^{N_{j}}\sim P_{S_{j}}$ denotes the $j$-th domain, and $P_{S_{i}}\neq P_{S_{j}}$ for $1\leq i,j\leq M$. $N_{j}$ is the number of total samples in $S_{j}$. The goal of domain generalization is to learn the predictive function $h$ in $M$ source domains to achieve minimum error on an unseen test domain $\mathcal{S}_{test}\sim P_{S_{test}}$, and $P_{S_{test}}\neq P_{S_{j}}$ for $1\leq j\leq M$:
$min~{}\mathbb{E}_{(X;Y)\in\mathcal{S}_{test}}\left[\mathcal{L}(h(X),Y)\right],$ (4)
where $X=(X_{f},X_{b},X_{a})$, $Y$ is the label, $\mathcal{L}$ is the loss function, and $\mathbb{E}$ is the expectation.
When $M=1$, it is a Single-to-single Cross-domain Generalization task, where the modal is trained on one training dataset and tested on another dataset. When $M\geqslant 2$, we propose three strategies to learn from multiple domains for the Multi-to-single Cross-domain Generalization. Let $B$ denote one batch of training data with a size of $N_{B}$. Given multiple training domains $\mathcal{S}_{train}=\{S_{j}|j=1,\cdots,M\}$, $B$ is a set of training data sampled from $\mathcal{S}_{train}$.
Domain-Simultaneous means to train multiple domains in parallel within each batch of data. In domain simultaneous training, the $k-$th batch of training data is a group of samples from different domains, i.e., $B^{k}=(b_{S_{1}}^{k},\cdots,b_{S_{M}}^{k})$, $k\in[1,\cdots K]$, where $b_{S_{j}}^{k}$ is the batch samples from domain $S_{j}$ for $j=1,\cdots,M$, $K$ is the number of batches during training. The total number of $b_{S_{j}}^{k}$ is $N_{B}$. As shown in Fig. 2 (b), each training batch contains smaller batch samples from all the source domains during training. Models are trained to learn from different domains simultaneously by feeding the mixed batch data.
[…]
3.3 Attention-Mixer Fusion
Besides investigating cross-domain sampling strategies, inspired by [19], we propose Attention-Mixer Fusion to enhance the performance by fusing audio-visual modalities, where the attention mixer layer takes multimodal features as input to produce fused features. In particular, an attention mixer layer is composed of unimodal MLP layers, self-attention layers [37], and crossmodal MLP layers. First, for batch size $N_{B}$, the input features from different modalities are concatenated and projected to be a tensor $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ by a Liner layer, followed by several attention mixer layers, where $N_{m}$ is the number of input modalities. Specifically, the unimodal MLP layer, the self-attention layer, and the crossmodal MLP layer can be respectively described as
$\small U^{*,*,i}=F_{g}^{*,*,i}+\mathbf{W_{2}}\ \sigma(\mathbf{W_{1}}\ LN(F_{g}^{*,*,i})),\ i=\left[1,D\right],$ (5)
$\footnotesize U=\left[\left(softmax\left(U{\mathbf{W_{3}}(U\mathbf{W_{4}})^{T}}\over{\sqrt{D}}\right)U\mathbf{W_{5}}\right)_{h}\right]\mathbf{W_{6}},\ h=[1,H],$ (6)
$\small U^{*,j,*}=U^{*,j,*}+\mathbf{W_{8}}\ \sigma(\mathbf{W_{7}}\ LN(U^{*,j,*})),\ j=\left[1,N_{m}\right],$ (7)
where $LN(\cdot)$ denotes the Layer Normalization, $\mathbf{W}_{1-8}$ are trainable weights, $H$ is the number of heads in multihead self-attention, and * denotes all the entries in that dimension. Several attention mixer layers are stacked as a deep block, which is set as a hyperparameter in practice. We set it to 6 in our experiment. Finally, $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ is reduced to $U\in\mathbb{R}^{N_{B}\times N_{m}\times 1}$ by obtaining the mean value on the feature dimension. In Eq. 5, the unimodal MLP layer is conducted along the feature dimension to learn the dynamics in each unimodal feature. Eq. 6 shows the multi-head self-attention operation on the tensor $U$, which further explores the attention between the unimodal features. In Eq. 7, the crossmodal MLP layer learns the dynamics across the modality dimension from the corresponding feature tokens.
TABLE I: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
Visual (AU) LSTM [38] 48.11 - - 61.21 - - - - - -
Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52
Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80
Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79
Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58
Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41
Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43
Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55
TABLE II: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
Domain-Simultaneous
Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 50.51
Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 52.53
Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 51.82
Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 55.04
Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.4 63.37 55.82
Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.5 51.12
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 51.78
Domain-Alternating
Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 53.50
Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 56.25
Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 54.25
Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 55.05
Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 55.66
Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 53.17
Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.57
Domain-by-Domain
Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 52.68
Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.46
Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 52.68
Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 55.05
Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 56.67
Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 54.24
Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 54.90
4 Experiments
In this part, extensive experiments are conducted to benchmark the cross-domain performances on public deception detection datasets. In the following, we sequentially describe the benchmark datasets & metrics (Sec. 4.1), implementation details (Sec. 4.2), benchmarking results (Sec. 4.3 - 4.4) and fusion performances (Sec. 4.5).
4.1 Databases and Metrics
Datasets. We benchmarked the cross-domain generalization performance based on four publicly available datasets. Real Life Trials [1] dataset is a popular real-world dataset collected from public court trials, which consists of 121 videos including 61 deceptive and 60 truthful video clips. As it is a real-world dataset, the Real Life trial dataset has more noise on both the video and audio. We filtered out some corrupted videos and obtained 108 videos (54 truthful and 54 deceptive) with 58 subjects for our experiments. Bag of Lies [2] is a multimodal dataset collected from well-controlled lab-based scenarios, where video, audio, EEG, and gaze data are collected. It has 35 subjects, 163 truthful and 162 deceptive video clips. The backgrounds for the videos are relatively clean and it is less noisy. MU3D [3] has 320 video clips and 80 subjects that cover different races and genders. It is also a lab-based dataset that uses the personal description paradigm to stimuli real-world cases. Each participant tells a positive truth, a positive lie, a negative truth, and a negative lie. Box of Lies [4] is a deception dataset collected from an online gameshow, which has 25 videos and 26 participants (6 male and 20 female). The full video set contains 29 truthful and 36 deceptive rounds of games. However, the quality of the original Box of Lies dataset is not satisfactory. The visual (the face of the participant) and audio from many clips are not matching due to the frequent changes of viewpoints. To perform a fair comparison, we preprocessed and cleaned the Box of Lies dataset. After preprocessing, 101 video clips were extracted for testing. Some of the typical samples from these datasets are shown in Fig. 1.
Evaluation Metrics. In this work, we followed the widely adopted metric, binary classification accuracy (%), for experimental evaluation. The deceptive clips were labeled as 1 and truthful clips were labeled as 0.
4.2 Implementation Details
Feature Extraction. Several widely-adopted audio and visual features were extracted by different tools. For visual features, OpenFace [32] was used to extract 35-dimensional AUs and 8-dimensional gaze features. Face frames were extracted and aligned by MTCNN [39], where we uniformly sampled 64 face frames for each video clip. Affect features were extracted by Emonet [33], where the feature included 5-class emotions, arousal, and valence. For audio features, Mel Spectrograms were extracted by OpenSmile toolkit [34]. Raw audio waveforms were also used in our experiments.
Protocols. Inspired by [35, 36], we treated each dataset as a domain. To evaluate the models’ cross-domain generalization capacity and alleviate domain information leakage, all the preprocessed data including original training and test data from each dataset was used for either training or testing. Note that the Box of Lies dataset was only used for testing as many samples were filtered out due to their unsatisfactory quality. The experiments were conducted on the single-to-single domain (e.g., R to B1 stands for training on Real-life Trial (R) and testing on Bag of Lies (B1)) and multi-to-single domain (e.g., R&M to B2 stands for training on Real-life Trial (R) and MU3D (M) and testing on Box of Lies (B2)).
Model Selection. Models for audio and visual modalities were selected to fit the data volume. For face frames, we adopted ResNet18 [31] and Gate Recurrent Unit (GRU) [40] models for facial feature extraction and temporal modeling, respectively. Two-layer multilayer perception (MLP) [41] models were used for AUs, gaze, and affect feature representation. For the audio-based Mel spectrogram, we used the ResNet18 [31] model for time-frequency feature representation. For audio waveforms, the Wave2Vec [42] model was applied for audio feature extraction.
Experimental Setting. Our proposed method was implemented with Pytorch. The ImageNet pretrained models (e.g., ResNet18) for classification were trained on the benchmark datasets using SGD optimizer with the initial learning rate (lr), momentum, and weight decay (wd) were 1e-3, 0.9, and 5e-5, respectively. We trained models with a maximum of 30 epochs and batchsize 32 on a single Nvidia V100 GPU. As for the fusion models (e.g., Atten-Mixer on face frames and Mel Spectrogram), Adam optimizer with initial lr=1e-3 and wd=5e-5 was used. The models were trained with batchsize 16 for a maximum of 30 epochs.
TABLE III: The fusion results of single-to-single cross-domain generalization accuracy (%).
Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
Score-level Average 53.23 41.58 51.88 64.49 62.38 50.62 57.94 53.23 62.38 55.30
Concat 51.08 55.45 51.88 54.21 58.42 51.25 57.01 50.15 61.39 54.54
SE-Concat 53.85 60.40 51.25 55.14 58.42 50.62 56.07 51.38 65.35 55.83
Cross-Atten 55.38 61.39 52.19 51.40 60.40 51.25 55.14 56.31 60.40 55.98
MLP-Mixer 55.08 48.51 53.44 56.07 58.42 53.75 55.14 59.08 59.41 55.43
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.92 59.41 57.94 63.37 53.75 53.75 60.75 56.00 61.39 58.14
Score-level Average 53.23 49.50 51.88 50.47 59.41 53.75 65.42 56.31 49.50 54.39
Concat 50.77 53.47 50.47 62.38 51.56 51.88 54.21 52.62 58.42 53.98
SE-Concat 50.15 44.55 51.40 61.39 53.12 52.50 65.42 56.31 66.34 55.69
Cross-Atten 54.46 51.49 55.14 58.42 51.25 52.19 63.55 56.62 66.34 55.95
MLP-Mixer 52.31 55.45 57.94 63.37 53.12 51.25 64.49 57.85 62.38 57.57
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.54 55.45 56.07 61.39 50.94 53.12 67.29 57.85 64.36 58.22
Score-level Average 49.85 58.42 54.06 45.79 60.40 50.62 49.53 56.00 63.37 54.23
Concat 49.54 53.47 47.66 61.39 53.44 51.88 57.01 50.15 58.42 53.66
SE-Concat 50.15 48.51 55.14 60.40 53.12 50.00 57.01 49.85 63.37 54.17
Cross-Atten 53.23 44.55 57.94 55.45 54.69 54.06 63.55 51.69 64.36 55.50
MLP-Mixer 49.54 57.43 50.47 63.37 54.06 53.12 59.81 55.08 69.31 56.91
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.52 54.46 57.94 61.39 53.12 51.56 69.16 58.15 64.36 58.18
Score-level Average 52.00 58.42 51.88 53.27 58.42 50.94 57.01 53.54 61.39 55.20
Concat 53.23 58.42 55.14 61.39 52.50 52.19 56.07 54.15 60.40 55.94
SE-Concat 51.38 58.42 51.40 62.38 52.81 50.94 59.81 54.77 52.48 54.93
Cross-Atten 51.08 48.51 55.14 60.40 53.12 53.44 60.75 56.31 60.40 55.46
MLP-Mixer 55.69 46.53 44.86 63.37 51.56 50.94 64.49 56.00 60.40 54.87
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 60.40 57.01 64.36 53.44 51.25 67.29 56.00 62.38 58.58
4.3 Cross-domain Testing with Unimodal Features
In this subsection, we present the benchmark results of cross-domain testing by investigating unimodal features to evaluate their generalization capacities. For clarity, we use “visual (face frames)” to indicate face inputs and “visual (AU/gaze/affect)” to indicate behavior inputs.
Single-to-Single Domain. Specifically, the models were trained on one dataset from one domain and tested on the other dataset from another domain. The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M). As shown in Tabel I, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect. For audio modality, Mel spectrogram and waveform were extracted. We also applied several different backbone networks as audio and visual encoders. We can observe that R and B1 datasets generalized the best on B2, and M generalized the best on R. Note that the B2 dataset was not adopted as a source domain dataset as the original dataset had too much noise and we cleaned it only for testing. On average, we can observe that the best result was achieved by using visual (AU+gaze+affect) features.
Multi-to-Single Domain. Here we fully evaluate the performance of multi-to-single cross-domain generalization using unimodal features. We conducted experiments for different domain sampling strategies. As shown in Table II, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%). For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%). The results showed that the best generalization performances were obtained when transferring to the real-world dataset and the gameshow dataset by training on the lab-based datasets. This was because the lab-based datasets (Bag of Lies and MU3D) are relatively clean compared to the real-world dataset (Real-life Trial) and the gameshow dataset (Box of Lies). However, in the opposite case, the generalization performance degraded, for example, R&M to B1 and R&B1 to M. Different domain sampling strategies reached their best average performance on different input features and backbone networks. To be specific, for both domain-simultaneous and domain-by-domain strategies, models trained on visual (AU+gaze+affect) features reached their highest accuracies, which were 55.82% and 56.67%, respectively. Using the domain-alternating strategy, the best accuracy of 56.25% was achieved by training on visual (face frames) features. We can observe that models trained on visual modalities outperformed those trained on audio modalities across all the generalization strategies. This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
Following the implementation by Ganin et al. [43], we compared the multi-to-single domain generalization accuracies w/w.o GRL. GRL was proposed to mitigate the domain shift issue by manipulating the training gradients. It worked by acting as an identity transform in forward propagation and multiplying the gradient by a certain negative constant during the backpropagation without having trainable parameters. GRL was inserted between encoders and domain classifiers, which was easy to implement. As GRL is a widely adopted method for domain generalization, it is investigated to show its effectiveness for the deception detection task. We selected domain-simultaneous as the baseline and added GRL to the original network with the same training setups. The average accuracies were reported in Fig. 3, where different types of visual and audio features and methods were compared. Training with GRL, the performance of ResNet18 and ResNet18+GRU models using visual (face frames) features and Wave2Vec model using waveform were enhanced. However, we observed that MLP models using visual (AU/gaze/affect) features and the ResNet18 model using Mel spectrograms degraded in performance. Generally, ResNet18 trained with GRL performed better than MLP for visual modality, and Wave2Vec trained with GRL boosted the performance and surpassed the model trained on the Mel spectrogram.
Figure 3: Performance comparisons of Domain-simultaneous training w/ and w/o Gradient Reversal Layer (GRL).
TABLE IV: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
Domain-Simultaneous
Score-level Average 53.23 43.56 57.43 50.94 51.40 58.42 48.51 51.93
Concat 52.00 56.44 60.40 51.88 51.40 62.38 57.43 55.99
SE-Concat 55.69 58.42 44.55 52.50 56.07 58.42 58.42 54.87
Cross-Atten 51.08 52.48 55.45 51.56 57.01 60.40 57.43 55.06
MLP-Mixer 53.85 43.56 62.38 52.19 56.07 61.39 59.41 55.55
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
Score-level Average 49.54 55.45 52.48 51.25 54.21 59.41 55.45 53.97
Concat 49.54 54.46 49.50 53.75 44.86 63.37 54.46 52.85
SE-Concat 52.00 58.42 42.57 51.56 54.21 58.42 63.37 54.36
Cross-Atten 50.46 52.48 60.40 53.12 52.34 58.42 59.41 55.23
MLP-Mixer 53.54 57.43 52.48 50.31 52.34 60.40 47.52 53.43
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
Score-level Average 48.62 58.42 56.44 50.94 51.40 56.44 53.47 53.68
Concat 48.31 53.47 56.44 50.31 56.07 63.37 57.43 55.06
SE-Concat 49.85 50.50 47.52 52.19 49.53 57.43 61.39 52.63
Cross-Atten 51.38 63.37 61.39 51.25 48.60 59.41 58.42 56.26
MLP-Mixer 55.69 59.41 57.43 53.75 50.47 61.39 50.50 55.52
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
Score-level Average 54.77 57.43 57.43 52.81 60.75 58.42 56.44 56.86
Concat 54.46 54.46 47.52 53.12 53.27 61.39 56.44 54.38
SE-Concat 53.23 55.45 54.46 52.50 59.81 56.44 58.42 55.76
Cross-Atten 48.62 57.43 50.50 53.12 48.60 58.42 56.44 53.30
MLP-Mixer 48.92 48.51 59.41 52.81 54.21 57.43 51.49 53.25
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
Domain-Alternating
Score-level Average 55.38 57.43 58.42 50.94 57.94 50.50 61.39 56.00
Concat 55.08 58.42 63.37 51.56 62.62 62.38 53.47 58.13
SE-Concat 50.46 49.50 63.37 52.19 57.01 65.35 56.44 56.33
Cross-Atten 55.08 59.41 60.40 52.50 48.60 58.42 61.39 56.54
MLP-Mixer 57.54 63.37 63.37 50.94 65.42 52.48 56.44 58.51
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.62 61.39 63.37 51.56 58.88 62.38 58.42 58.95
Score-level Average 50.15 58.42 60.40 50.62 51.40 60.40 58.42 55.69
Concat 53.23 62.38 54.46 51.25 59.81 54.46 54.46 55.72
SE-Concat 50.15 52.48 64.36 50.31 59.81 55.45 50.50 54.72
Cross-Atten 52.92 57.43 49.50 52.81 65.42 53.47 66.34 56.84
MLP-Mixer 51.69 58.42 57.43 51.25 62.62 58.42 56.44 56.61
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.23 63.37 56.44 51.25 63.55 59.41 56.44 57.67
Score-level Average 51.38 56.44 54.46 50.00 64.49 61.39 49.50 55.38
Concat 49.85 62.38 58.42 51.88 61.68 55.45 51.49 55.88
SE-Concat 50.46 62.38 58.42 50.00 57.94 59.41 65.35 57.71
Cross-Atten 49.85 62.38 50.50 53.44 58.88 56.44 56.44 55.42
MLP-Mixer 50.15 55.45 58.42 50.31 58.88 61.39 57.43 56.00
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 50.15 59.41 50.50 52.19 58.88 60.40 61.39 56.13
Score-level Average 56.00 52.48 59.41 51.88 51.40 60.40 56.44 55.43
Concat 51.69 58.42 58.42 52.81 57.01 61.39 56.44 56.60
SE-Concat 56.31 59.41 58.42 50.31 53.27 61.39 55.45 56.37
Cross-Atten 49.54 67.33 54.46 51.56 66.36 52.48 60.40 57.45
MLP-Mixer 50.15 63.37 60.40 51.56 62.62 61.39 61.39 58.70
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 66.34 60.40 51.88 59.81 61.39 62.38 59.13
Domain-by-Domain
Score-level Average 54.77 46.53 59.41 50.62 57.94 59.41 58.42 55.30
Concat 53.85 53.47 63.37 51.88 57.94 65.35 58.42 57.75
SE-Concat 56.62 44.55 63.37 52.19 54.21 58.42 59.41 55.54
Cross-Atten 54.15 49.50 62.38 54.69 57.01 60.40 63.37 57.36
MLP-Mixer 57.54 62.38 71.29 51.25 54.21 62.38 62.38 60.20
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 57.54 58.42 63.37 52.19 69.16 60.40 64.36 60.78
Score-level Average 52.00 50.50 58.42 51.56 57.94 61.39 60.40 56.03
Concat 50.46 59.41 62.38 52.19 60.75 56.44 57.43 57.01
SE-Concat 55.38 48.51 58.42 52.81 62.62 57.43 59.41 56.37
Cross-Atten 52.31 53.47 58.42 51.25 64.49 58.42 57.43 56.54
MLP-Mixer 53.23 60.40 62.38 52.19 61.68 58.42 58.42 58.10
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.85 66.34 63.37 53.75 58.88 64.36 63.37 60.56
Score-level Average 48.62 59.41 51.49 51.25 60.75 62.38 63.37 56.75
Concat 57.23 60.40 56.44 51.25 56.07 57.43 56.44 56.47
SE-Concat 55.08 60.40 59.41 54.06 60.75 61.39 57.43 58.36
Cross-Atten 50.15 56.44 61.49 57.50 59.81 64.36 58.42 58.31
MLP-Mixer 52.62 56.44 58.42 52.50 60.75 61.39 58.42 57.22
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 56.44 57.43 54.06 66.36 64.36 60.40 58.68
Score-level Average 53.85 51.49 65.35 53.75 58.88 57.43 58.42 57.02
Concat 52.92 52.48 55.45 53.75 62.62 62.38 56.44 56.58
SE-Concat 51.38 60.40 58.42 51.56 57.01 62.38 59.41 57.22
Cross-Atten 51.80 52.48 62.38 53.12 59.81 64.36 60.40 57.76
MLP-Mixer 53.85 55.45 60.40 50.94 59.81 64.36 60.40 57.89
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 61.39 59.41 55.00 58.88 59.41 60.40 58.60
4.5 Cross-domain Testing with Multimodal Fusion
Here we present multimodal fusion results of cross-domain testing to evaluate models’ generalization capacities.
Single-to-Single Domain with Fusion. In this section, we conducted experiments of single-to-single cross-domain testing. Two types of fusion positions were involved including score-level and feature-level fusion. For feature-level fusion, multiple fusion methods were adopted such as simple concatenation, SE-Concat [44], Cross-Atten [37], MLP-Mixer [19], and Attention-Mixer fusion (Ours). To be specific, simple concatenation refers to the concatenation of the extracted features before the input to the classifier. SE-concat stands for the SE attention applied to the concatenated features. Cross-Atten means the crossmodal attention among input features by using the attention mechanism from the Transformer. MLP-Mixer uses the method in  [19] on the extracted features. The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram). It results in four combinations of either two or three types of inputs. As shown in Table IV, among each type of input combination, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results. The rest fusion methods showed comparable results, which were less performed than Atten-Mixer.
Multi-to-Single Domain with Fusion. We benchmarked multi-to-single cross-domain generalization with different fusion methods by three types of cross-testing strategies. In total, seven sub-experiments were conducted for different domain combinations. The results are shown in Tabel IV. On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively. Among these, for both domain-simultaneous and domain-by-domain strategies, the best results were achieved by taking visual (face frames) and visual (AU+gaze+affect) features as input, while the best result for domain-alternating was achieved by using visual (face frames), visual (AU+gaze+affect) and audio (Mel spectrogram) features. Taking a close look at the average fusion results, for each type of modality input, Atten-Mixer achieved the best among the six fusion methods. showing the effectiveness of the proposed method. By using Atten-Mixer, in general, the average result showed slightly better when using the domain-by-domain strategy and taking visual (face frames) and visual (AU+gaze+affect) features as input. To sum up, the results showed that on the current publicly available datasets, visual features were better when it came to multi-to-single cross-domain generalizability. However, the performance differences were not a large gap, and there were also no significant differences by comparing two-to-one domain and three-to-one domain cross-testing performances.
Figure 4: Ablation study for attention-mixer layers. The number of layers 4, 5, 6, and 7 are compared. The modality and inputs for A, B, C, and D are in line with those in Table III from the top to the bottom.
Ablation Study for Attention-Mixer Fusion Module. We conducted an ablation study for the proposed attention-mixer fusion module with the changes in the number of attention-mixer layers. The experiments were conducted on single-to-single domain testing, where the average accuracies were compared. As shown in Fig. 4, the number of attention-mixer layers was set to 4, 5, 6, and 7. The modalities and inputs in Table III are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram). The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
4.6 Discussion
We can observe that the general performance of cross-domain deception detection is unsatisfactory because it is challenging to reduce the domain gap between each dataset. The domain generalization ability of widely-adopted methods was relatively weak using either audio or visual features. Different domain sampling strategies worked well for different audiovisual features. Fusing multiple modalities is able to mitigate the problem. However, the performance still needs to be improved.
Ethical Consideration. Developing deception detection using AI should emphasize respecting privacy, minimizing psychological harm, preventing discrimination, promoting transparency, etc. Researchers should follow appropriate regulations to develop and deploy AI systems for deception detection. Potential misuses and negative impacts include invasion of privacy, discrimination, erosion of trust, etc. Mitigating these risks requires responsible practices from researchers and developers.
5 Conclusion
In this paper, we benchmark the cross-domain generalization performance for deception detection on publicly available datasets. We compare the single-to-single domain and multi-to-single domain generalization performances, where three strategies are used including domain-simultaneous, domain-alternating, and domain-by-domain. We also investigate the effectiveness of the gradient reversal layer for domain-simultaneous strategy. Moreover, we propose the Attention-Mixer fusion method to alleviate the domain shift issue and boost the performance. Future works for deception detection are encouraged to propose better methods to improve the domain generalizability on audio-visual deception detection task.
Acknowledgments.  This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, Nanyang Technological University, Singapore. The research is supported by the DSO National Laboratories, under project agreement No. DSOCL21238.

### v2

@@
 To further exploit the impacts using data from multiple source domains for training, we investigate three types of domain sampling strategies, including domain-simultaneous, domain-alternating, and domain-by-domain for multi-to-single domain generalization evaluation.
+We also propose an algorithm to enhance the generalization performance by maximizing the gradient inner products between modality encoders, named “MM-IDGM”.
 Furthermore, we proposed the Attention-Mixer fusion method to improve performance, and we believe that this new cross-domain benchmark will facilitate future research in audio-visual deception detection.
@@
 Traditional deception detection is often a contact-based method.
-It assesses whether someone is telling the truth or not by monitoring physiological responses like skin conductance and heart rate [12].
+It assesses whether someone is telling the truth or not by monitoring physiological responses like skin conductance and heart rate [12, 13, 14].
 Experts’ Behavioral observation and analysis are another technique that evaluates changes in a person’s body language, speech patterns, and eye movements [15, 16].
@@
 Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing.
-To further enhance performance, we propose an Attention-Mixer fusion method based on MLP-Mixer [19].
+To further enhance performance, we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM) to improve multi-to-single domain generalization performance and an Attention-Mixer fusion method based on MLP-Mixer [21].
+In our experiments, we compared single-to-single and multi-to-single domain generalization, incorporating three sampling strategies, one domain generalization method, and six fusion methods across two modalities.
+The benchmarking results among different features reveal that the best average fusion accuracy was 58.58% for single-to-single cross-domain generalization and 59.13% for multi-to-single cross-domain generalization.
+Moreover, for MM-IDGM, we achieved the highest average accuracy of 59.69% based on domain-simultaneous sampling.
 This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings.
@@
 In summary, our main contributions include:
-  • Introducing a new benchmark for evaluating the generalization capacity of AI models using audio and visual features across different domains.
-  • Comparing the single-to-single domain and multi-to-single domain generalization using different architectures.
-  • Providing three domain sampling strategies i.e., domain simultaneous, domain alternating, and domain-by-domain, to conduct multi-to-single cross-domain testing.
-  • Proposing the Attention-Mixer fusion method to enhance performance.
+  • Introducing a new benchmark for cross-domain audio-visual deception detection: We present a comprehensive benchmark that evaluates the generalization capacity of AI models using audio and visual modalities across multiple domains in the deception detection task.
+  • Comparing single-to-single and multi-to-single domain generalization across different architectures: We conduct extensive experiments comparing the performance of AI models in single-to-single and multi-to-single domain generalization over different architectures and modalities inputs.
+  • Testing three domain sampling strategies for multi-to-single domain generalization: We propose and evaluate three distinct domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, to assess their effectiveness and offer diverse approaches to handling domain discrepancies.
+  • Introducing Multimodal Inter-Domain Gradient Matching (MM-IDGM) for domain generalization: We propose MM-IDGM to improve multi-to-single domain generalization by aligning gradient updates between modalities, enhancing the model’s ability to transfer knowledge across different domains.
+  • Proposing the Attention-Mixer fusion method for enhanced multimodal performance: We introduce the Attention-Mixer fusion method that combines MLP-Mixer and self-attention layers to capture both intra- and inter-modal interactions better, leading to improved fusion performances in audio-visual deception detection tasks.
 In the rest of the paper, Sec.
@@
 They argued that different psychological states can be exploited by adopting appropriate interview techniques of liars and truth-tellers.
-Hirschberg et al.
-[26] proposed a method to distinguish deceptive from non-deceptive speeches by a large corpus.
-They also conducted experiments using acoustic, lexical, and speaker-dependent features, which showed improved performance by combining multiple feature sets.
 Warren et al.
 [28] conducted experiments to investigate the relationship between affective facial expressions to deception.
-The results indicated that leaked emotions with the incongruous intended message can provide useful cues to deception, which supported the nonverbal leakage theory [28, 29].
+The results indicated that leaked emotions with the incongruous intended message can provide useful cues to deception, which supported the nonverbal leakage theory [29].
 Figure 2: Main network and method.
@@
 2.2 Multimodal Deception Detection
-Recent works for deception detection usually use verbal and non-verbal features and propose effective fusion methods [30, 15, 16, 17].
-For example, some works utilized facial features from RGB images to perform deception detection [15, 17, 5].
-To capture facial movements, facial action units (AUs) were utilized.
-Other features, such as facial expression, were also adopted [6, 30].
-Besides visual features, many works incorporated audio features to boost performance [7, 6, 15].
+Recent advances in deception detection have increasingly incorporated both verbal and non-verbal cues, leveraging multimodal data to enhance detection performance.
+Often, those methods integrate features from different modalities, such as visual, audio, and textual modalities, and propose fusion strategies to combine them efficiently [30, 17, 18, 19].
+In this section, we review and discuss the strengths and limitations of recent audio-visual deception detection methods, including the fusion techniques used and the gaps in addressing cross-domain issues.
+Visual Features in Deception Detection.
+Many recent works have utilized facial features extracted from RGB images to perform deception detection [17, 19, 5].
+They focus on capturing facial movements and expressions, which are crucial indicators of deception.
+Many studies have employed Facial Action Units (AUs) as a key feature [17, 19] to analyze subtle facial movements.
+AUs provide detailed representations of specific facial muscle activations, making them highly useful for identifying deceptive cues.
+Additionally, gaze movements and facial expressions have been widely adopted as visual features in deception detection models [6, 30, 31, 32].
+By capturing emotional states that may signal deception, these methods can further enhance the detection capabilities.
+Audio Features for Deception Detection.
+Along with visual cues, many recent works have incorporated audio features to boost the performance of deception detection systems.
+Audio signals can provide complementary information, such as changes in speech patterns, tone, and hesitation, which are often associated with deceptive behavior [7, 6, 17].
 For example, Wu et al.
-[6] used MFCC (Mel-frequency Cepstral Coefficients) features and Karimi et al.
-[15] used raw audio.
-Most of the recent works mentioned above have considered multimodal fusion approaches that extract visual, audio, and text information to boost performance.
-In addtion to visual, audio and text, Karnati et al.
+[6] used MFCC (Mel-frequency Cepstral Coefficients) features, and Karimi et al.
+[17] explored the use of raw audio signals for deception detection, demonstrating the potential of unprocessed data to capture subtle vocal cues indicative of deception.
+Multimodal Fusion Techniques.
+The integration of multiple modalities, such as visual, audio, and textual information, has been shown to improve deception detection performance significantly.
+Most recent works mentioned above have considered multimodal fusion approaches that extract visual, audio, and text information to boost performance [33, 34].
+In addition to visual, audio, and text, Karnati et al.
 [18] exploited physiological signals, i.e., EEG representations for deception detection.
-To better fuse the multimodal features, different fusion methods were proposed, which can be broadly categorized into feature-level fusion and decision-level fusion.
-Specifically, feature-level fusion focused on producing better multimodal embeddings and used the linear layers to extract crossmodal dynamics [15, 17, 5, 6, 7].
-Rather than that, decision-level fusion aimed to fuse multimodal dynamics at a late stage, to reduce computational complexity and learn good marginal representations [16, 7].
-However, previous works on multimodal deception detection did not consider cross-domain issues that occur from one domain to another, which is the focus of this work.
+Fusion methods can be broadly categorized into two types: feature-level fusion and decision-level fusion.
+Feature-level fusion combines raw or intermediate features from each modality into a single representation, allowing the model to learn cross-modal dynamics [17, 19, 5, 6, 7].
+This method often uses different encoders, linear layers, or fusion modules to blend features and capture interactions between modalities.
+On the other hand, decision-level fusion merges the outputs or produced logits of individual unimodal models at a later stage.
+It can reduce computational complexity and enable the system to learn effective marginal representations from each modality [18, 7].
+Both approaches have shown promising results in improving the performance of deception detection systems.
+Cross-Domain Generalization Despite significant progress, there still exists the challenge of cross-domain generalization in multimodal deception detection.
+Previous works have mainly focused on optimizing unimodal or fusion methods within a single domain, without considering the domain shift when the system is applied to different environments or populations.
+For example, models trained on controlled lab data may not generalize well to real-world settings, where the recording conditions and communication styles may introduce significant variability.
+In this work, we aim to build a benchmark for cross-domain generalization performance on the widely used audio-visual deception detection datasets, which is crucial for evaluating and improving the robustness of deception detection models in different scenarios.
+Establishing such a benchmark will provide clearer comparisons, highlight model weaknesses, and guide the development of systems across different domains.
 3 Methodology
@@
 As depicted in Fig.
-1, this network structure offers several advantages: (1) flexibility in network selection: different audio or visual encoders can be effortlessly incorporated and compared in a fair manner, (2) adaptability: the addition or removal of specific modules and/or losses is straightforward.
+2, this network structure offers several advantages: (1) flexibility in network selection: different audio or visual encoders can be effortlessly incorporated and compared in a fair manner, (2) adaptability: the addition or removal of specific modules and/or losses is straightforward.
 For instance, a fusion module can be inserted before classifiers, and (3) easy performance benchmarking: the system facilitates evaluating performance in various settings, such as score-level fusion and feature-level fusion.
@@
 As shown in Fig.
-1, given a detected RGB face image as input $X_{f}$, the deep features $F_{f}$ could be extracted via face encoder networks $\mathcal{E}_{f}$ (e.g., ResNet18 [31]).
+2, given a detected RGB face image as input $X_{f}$, the deep features $F_{f}$ could be extracted via face encoder networks $\mathcal{E}_{f}$ (e.g., ResNet18 [35]).
 Similarly, behavior inputs such as the AU and/or affect features $X_{b}$ are encoded by OpenFace [36] or affect model (e.g., EmotionNet [37]) $\mathcal{E}_{b}$ to output behavior features $F_{b}$.
@@
 The overall loss function can be described as follows:
-$\mathcal{L}=\frac{1}{N}\sum_{i=i}^{N}\left(\sum_{m={f,b,a}}\mathcal{L}_{m,i}+\lambda\mathcal{L}_{g,i}\right),$ (3)
-where $N$ is the number of data samples and $\lambda$ is a trade-off parameter between modality loss and fusion loss. $\lambda$ is set to 0.5 in our experiments.
+$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{m={f,b,a}}\mathcal{L}_{m,i}+\lambda\mathcal{L}_{g,i}\right),$ (3)
+where $N$ is the number of data samples and $\lambda$ is a trade-off parameter between modality loss and fusion loss.
+$\lambda$ is set to 0.5 in our experiments.
+For score fusion, the fusion score is the average of the individual predicted scores for each class, and the cross-entropy loss function is applied to the fused score.
+For feature fusion, individual features are the inputs for the fusion block.
 3.2 Cross-domain Generalization
@@
 $min~{}\mathbb{E}_{(X;Y)\in\mathcal{S}_{test}}\left[\mathcal{L}(h(X),Y)\right],$ (4)
-where $X=(X_{f},X_{b},X_{a})$, $Y$ is the label, $\mathcal{L}$ is the loss function, and $\mathbb{E}$ is the expectation.
+where $X=(X_{f},X_{b},X_{a})$ is the input, $Y$ is the label, $\mathcal{L}$ is the loss function, and $\mathbb{E}$ denotes the expectation.
 When $M=1$, it is a Single-to-single Cross-domain Generalization task, where the modal is trained on one training dataset and tested on another dataset.
 When $M\geqslant 2$, we propose three strategies to learn from multiple domains for the Multi-to-single Cross-domain Generalization.
+Previous works [41, 42] have indicated that different sampling strategies for domain generalization may result in different performances.
+Therefore, in this work, we propose three multi-to-single domain generalization sampling methods and compare their performance.
 Let $B$ denote one batch of training data with a size of $N_{B}$.
@@
 Domain-Simultaneous means to train multiple domains in parallel within each batch of data.
-In domain simultaneous training, the $k-$th batch of training data is a group of samples from different domains, i.e., $B^{k}=(b_{S_{1}}^{k},\cdots,b_{S_{M}}^{k})$, $k\in[1,\cdots K]$, where $b_{S_{j}}^{k}$ is the batch samples from domain $S_{j}$ for $j=1,\cdots,M$, $K$ is the number of batches during training.
+In domain simultaneous training, the $k-$th batch of training data is a group of samples from different domains, i.e., $B^{k}=(b_{S_{1}}^{k},\cdots,b_{S_{M}}^{k})$, $k\in[1,\cdots K]$, where $b_{S_{j}}^{k}$ is the batch samples from domain $S_{j}$ for $j=1,\cdots,M$ and $K$ is the number of batches during training.
 The total number of $b_{S_{j}}^{k}$ is $N_{B}$.
@@
 Models are trained to learn from different domains simultaneously by feeding the mixed batch data.
+This approach helps the model learn domain-invariant features by exposing it to data from multiple domains simultaneously.
+However, it may cause the model to focus on “average” features across domains, leading to poor generalization to the unseen domains.
+Mixing domains ensures broader feature exposure, but can dilute domain-specific nuances that are useful for recognizing patterns in unseen domains.
 […]
 […]
+Multimodal Inter-Domain Gradient Matching.
+Beyond the benchmarking on the three different multi-to-single sampling strategies, inspired by Shi et al.
+[43], we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM).
+IDGM was proposed by Shi et al.
+[43] and uses an effective optimization algorithm named “Fish” to approximate the second-order derivative in IDGM and to reduce computational cost.
+The main idea of IDGM is to maximize the gradient inner product to align the gradient directions across domains, so as to learn the weights that can produce the closer input-output correspondence.
+However, considering the multimodal nature of our task, directly optimizing IDGM in multimodal models may not be desirable.
+Different from IDGM in [43] which aims to learn the features that are invariant across domains, MM-IDGM prioritizes learning the invariant features across domains for the same modalities by maximizing the gradient inner product between modality encoders.
+These inner gradient products are then dynamically adjusted based on the unimodal losses from previous steps.
+The remaining parts of the models are then optimized across domain by IDGM.
+The MM-IDGM algorithm is developed based on the “Fish” algorithm and is presented in Algorithm 1.
+Algorithm 1 MM-IDGM for Modalities $m_{1}$ and $m_{2}$
+$L_{m_{1}}$, $L_{m_{2}}$triangleright from the previous step
+while  iterations $\leq$ maximum iterations do
+     $\tilde{\theta}_{m_{1}}\leftarrow\theta_{m_{1}}$, $\tilde{\theta}_{m_{2}}\leftarrow\theta_{m_{2}}$, $\tilde{\theta}_{r}\leftarrow\theta_{r}$
+     for $S_{i}\in$permute$\{S_{1},S_{2},\cdots,S_{M}\}$ do
+         Sample batch $s_{i}\sim S_{i}$
+         $g_{m_{1},i}=\mathbb{E}_{s_{i}}[\frac{\partial L(X_{m_{1}},Y);\tilde{\theta}_{m_{1}}}{\partial\tilde{\theta}_{m_{1}}}]$triangleright encoder $m_{1}$
+         $g_{m_{2},i}=\mathbb{E}_{s_{i}}[\frac{\partial L(X_{m_{2}},Y);\tilde{\theta}_{m_{2}}}{\partial\tilde{\theta}_{m_{2}}}]$triangleright encoder $m_{2}$
+         $g_{r,i}=\mathbb{E}_{s_{i}}[\frac{\partial L(X_{r},Y);\tilde{\theta}_{r}}{\partial\tilde{\theta}_{r}}]$triangleright other
+         Update $\tilde{\theta}_{m_{1}}\leftarrow\tilde{\theta}_{m_{1}}-\alpha(\frac{L_{m_{2}}}{L_{m_{1}}+L_{m_{2}}})g_{m_{1},i}$
+         Update $\tilde{\theta}_{m_{2}}\leftarrow\tilde{\theta}_{m_{2}}-\alpha(\frac{L_{m_{1}}}{L_{m_{1}}+L_{m_{2}}})g_{m_{2},i}$
+         Update $\tilde{\theta}_{r}\leftarrow\tilde{\theta}_{r}-\alpha g_{r,i}$
+     end for
+     Update $\theta\leftarrow\theta+\epsilon(\tilde{\theta}-\theta))$
+     Update $L_{m_{1}}$, $L_{m_{2}}$
+end while
+MM-IDGM first performs inner gradient updates on a cloned version of the original model, denoted as $\tilde{\theta}$, focusing on each modality encoder individually.
+The gradients for each modality encoder are dynamically adjusted based on previous unimodal losses, ensuring a more balanced learning rate across modalities.
+Finally, the weights of the original model, $\theta$, are updated by applying a weighted difference between the cloned model and the original one.
 3.3 Attention-Mixer Fusion
@@
 In particular, an attention mixer layer is composed of unimodal MLP layers, self-attention layers [44], and crossmodal MLP layers.
-First, for batch size $N_{B}$, the input features from different modalities are concatenated and projected to be a tensor $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ by a Liner layer, followed by several attention mixer layers, where $N_{m}$ is the number of input modalities.
+First, for batch size $N_{B}$, the input features from different modalities are concatenated and projected to be a tensor $F_{g}\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ by a Liner layer, followed by several attention mixer layers, where $N_{m}$ is the number of input modalities.
 Specifically, the unimodal MLP layer, the self-attention layer, and the crossmodal MLP layer can be respectively described as
-$\small U^{*,*,i}=F_{g}^{*,*,i}+\mathbf{W_{2}}\ \sigma(\mathbf{W_{1}}\ LN(F_{g}^{*,*,i})),\ i=\left[1,D\right],$ (5)
-$\footnotesize U=\left[\left(softmax\left(U{\mathbf{W_{3}}(U\mathbf{W_{4}})^{T}}\over{\sqrt{D}}\right)U\mathbf{W_{5}}\right)_{h}\right]\mathbf{W_{6}},\ h=[1,H],$ (6)
-$\small U^{*,j,*}=U^{*,j,*}+\mathbf{W_{8}}\ \sigma(\mathbf{W_{7}}\ LN(U^{*,j,*})),\ j=\left[1,N_{m}\right],$ (7)
+$U^{*,*,i}=F_{g}^{*,*,i}+\mathbf{W_{2}}\ \sigma(\mathbf{W_{1}}\ LN(F_{g}^{*,*,i})),\ i=\left[1,D\right],$ (5)
+$\small U=\left[\left(softmax\left(U{\mathbf{W_{3}}(U\mathbf{W_{4}})^{T}}\over{\sqrt{D}}\right)U\mathbf{W_{5}}\right)_{h}\right]\mathbf{W_{6}},\ h=[1,H],$ (6)
+$U^{*,j,*}=U^{*,j,*}+\mathbf{W_{8}}\ \sigma(\mathbf{W_{7}}\ LN(U^{*,j,*})),\ j=\left[1,N_{m}\right],$ (7)
 where $LN(\cdot)$ denotes the Layer Normalization, $\mathbf{W}_{1-8}$ are trainable weights, $H$ is the number of heads in multihead self-attention, and * denotes all the entries in that dimension.
@@
 We set it to 6 in our experiment.
-Finally, $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ is reduced to $U\in\mathbb{R}^{N_{B}\times N_{m}\times 1}$ by obtaining the mean value on the feature dimension.
+Finally, the output tensor $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ is reduced to $U\in\mathbb{R}^{N_{B}\times N_{m}\times 1}$ by obtaining the mean value on the feature dimension.
 In Eq.
@@
 7, the crossmodal MLP layer learns the dynamics across the modality dimension from the corresponding feature tokens.
-TABLE I: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
-Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
-Visual (AU) LSTM [38] 48.11 - - 61.21 - - - - - -
-Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52
-Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80
-Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79
-Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58
-Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41
-Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43
-Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55
-TABLE II: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
-Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
-Domain-Simultaneous
-Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 50.51
-Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 52.53
-Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 51.82
-Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 55.04
-Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.4 63.37 55.82
-Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.5 51.12
-Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 51.78
-Domain-Alternating
-Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 53.50
-Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 56.25
-Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 54.25
-Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 55.05
-Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 55.66
-Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 53.17
-Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.57
-Domain-by-Domain
-Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 52.68
-Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.46
-Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 52.68
-Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 55.05
-Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 56.67
-Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 54.24
-Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 54.90
+In contrast to the MLP-Mixer, which utilizes a simple two-layer MLP structure, our approach introduces a self-attention layer between the two MLP layers.
+This allows for more efficient global operation from MLP-Mixer and also adaptively weighing token interaction by self-attention.
+This architectural enhancement allows our model to more effectively capture both intra- and inter-modal interactions, making it especially well-suited for processing multimodal data such as audio and visual inputs.
 4 Experiments
-In this part, extensive experiments are conducted to benchmark the cross-domain performances on public deception detection datasets.
-In the following, we sequentially describe the benchmark datasets & metrics (Sec.
-4.1), implementation details (Sec.
-4.2), benchmarking results (Sec.
-4.3 - 4.4) and fusion performances (Sec.
-4.5).
+In this part, extensive experiments are conducted to benchmark the cross-domain performances of public deception detection datasets.
 4.1 Databases and Metrics
@@
 Two-layer multilayer perception (MLP) [47] models were used for AUs, gaze, and affect feature representation.
+Specifically, the first MLP Layer initially operates on the dimension that represents the number of features, with sizes following the sequence: “input dimension $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 1.
+Afterward, it reshapes the tensor, and the MLP Layer processes on the feature dimension, where the sizes progress as “64 $\rightarrow$ 32 $\rightarrow$ 16 $\rightarrow$ 2.
+The MLP networks have identical structures, except for their first linear layers, as the input dimensions differ.
 For the audio-based Mel spectrogram, we used the ResNet18 [35] model for time-frequency feature representation.
@@
 The models were trained with batchsize 16 for a maximum of 30 epochs.
-TABLE III: The fusion results of single-to-single cross-domain generalization accuracy (%).
+TABLE I: The results of intra-domain testing accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), and MU3D (M).
+Modality & Inputs Method R to R B1 to B1 M to M
+Visual (Face frames) ResNet18 70.10 57.78 55.15
+Visual (AU+Gaze+Affect) MLP 62.87 55.84 55.23
+Audio (Mel spectrogram) ResNet18 75.78 55.69 56.45
+Visual (Face frames) concat 59.09 54.46 56.30
++ Visual (AU+Gaze+Affect) Atten-Mixer (Ours) 63.05 55.45 57.59
+Visual (Face frames) concat 68.75 54.74 56.19
++ Audio (Mel spectrogram) Atten-Mixer (Ours) 71.00 56.44 57.14
+Visual (AU+Gaze+Affect) concat 68.88 56.25 53.27
++ Audio (Mel spectrogram) Atten-Mixer (Ours) 69.30 56.78 55.23
+Visual (Face frames) concat 69.33 58.43 55.24
++ Visual (AU+Gaze+Affect) Atten-Mixer(Ours) 70.52 59.41 56.88
++ Audio (Mel spectrogram)
+TABLE II: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
+Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
+Visual (AU) LSTM 48.11 - - 61.21 - - - - - -
+Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52
+Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80
+Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79
+Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58
+Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41
+Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43
+Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55
+TABLE III: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
+Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
+Domain-Simultaneous
+Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 50.51
+Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 52.53
+Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 51.82
+Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 55.04
+Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.4 63.37 55.82
+Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.5 51.12
+Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 51.78
+Domain-Alternating
+Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 53.50
+Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 56.25
+Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 54.25
+Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 55.05
+Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 55.66
+Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 53.17
+Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.57
+Domain-by-Domain
+Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 52.68
+Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.46
+Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 52.68
+Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 55.05
+Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 56.67
+Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 54.24
+Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 54.90
+TABLE IV: The fusion results of single-to-single cross-domain generalization accuracy (%).
 Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
@@
 For clarity, we use “visual (face frames)” to indicate face inputs and “visual (AU/gaze/affect)” to indicate behavior inputs.
+Intra-Domain Testing.
+Before performing cross-domain testing, we show the results of intra-domain testing on Real-life Trial (R), Bag of Lies (B1), and MU3D (M) in Table I.
+It is important to note that we did not use the Box of Lies (B2) dataset for training due to concerns about its quality.
+Specifically, we performed 5-fold intra-domain testing and reported the average accuracies.
+ResNet18 and MLP were used as baseline models, as they are widely used as encoders in audio-visual deception detection.
+As expected, the performance of intra-domain testing varies across datasets, reflecting the differences in their levels of difficulty.
 Single-to-Single Domain.
@@
 The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
-As shown in Tabel I, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect.
+As shown in Tabel II, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect.
 For audio modality, Mel spectrogram and waveform were extracted.
@@
 We conducted experiments for different domain sampling strategies.
-As shown in Table II, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%).
+As shown in Table III, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%).
 For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%).
@@
 This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
+Discussion: (1) Wave2Vec outperforms MFCC in audio feature extraction due to its ability to capture richer, more meaningful representations, as demonstrated across various sampling methods.
+Additionally, larger and more diverse datasets like B1 and M enhance generalization, leading to improved performance when models trained on these datasets are tested on smaller or less challenging datasets like R.
+The R dataset, composed of courtroom trial videos, tends to yield better results, while B2, a game show dataset, often produces unsatisfactory outcomes due to its lower video quality and data imbalance.
+Overall, visual features perform better than audio features and models using Wave2Vec features achieve better results compared to those trained on MFCC.
+(2) While domain-simultaneous sampling is typically expected to enhance generalization by promoting domain-invariant learning, our average results show that domain-by-domain sampling slightly outperforms it in these audio-visual deception detection tasks.
+This may be due to its ability to capture domain-specific variations, which are beneficial in this context.
+However, the effectiveness of each sampling strategy depends on the features being used; for example, Wave2Vec features perform better with domain-simultaneous sampling, while MFCC features benefit more from domain-by-domain sampling.
+This suggests that the optimal sampling strategy may vary depending on the feature set’s predictive power and the specific dataset characteristics.
+(3) Visual features, particularly behavioral features like AUs, gaze, and affective cues, outperform audio features in most of our experiments.
+Behavioral features are more predictive in the majority of testing pairs, as seen in Table II, and achieved the best average performance across sampling strategies in Table III.
+Their rich information content, capturing subtle signs of deception, may explain their superior performance over the face or audio features.
+Due to the lower dimension compared with face and audio features, behavioral features make them more effective when training on small datasets for detecting deception.
 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
@@
 Figure 3: Performance comparisons of Domain-simultaneous training w/ and w/o Gradient Reversal Layer (GRL).
-TABLE IV: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
+TABLE V: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
 Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
@@
 4.5 Cross-domain Testing with Multimodal Fusion
-Here we present multimodal fusion results of cross-domain testing to evaluate models’ generalization capacities.
 Single-to-Single Domain with Fusion.
@@
 It results in four combinations of either two or three types of inputs.
-As shown in Table IV, among each type of input combination, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results.
-The rest fusion methods showed comparable results, which were less performed than Atten-Mixer.
+As shown in Table IV, among all the fusion methods, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results on average.
 Multi-to-Single Domain with Fusion.
@@
 In total, seven sub-experiments were conducted for different domain combinations.
-The results are shown in Tabel IV.
+The results are shown in Tabel V.
 On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively.
@@
 The number of layers 4, 5, 6, and 7 are compared.
-The modality and inputs for A, B, C, and D are in line with those in Table III from the top to the bottom.
+The modality and inputs for A, B, C, and D are in line with those in Table IV from the top to the bottom.
 Ablation Study for Attention-Mixer Fusion Module.
@@
 4, the number of attention-mixer layers was set to 4, 5, 6, and 7.
-The modalities and inputs in Table III are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram).
+The modalities and inputs in Table IV are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram).
 The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
+TABLE VI: The fusion results of multi-to-single cross-domain generalization accuracy (%) for multimodal inter-domain gradient matching (MM-IDGM).
+The numbers in the brackets show the improved accuracies compared with Atten-Mixer.
+Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
+Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
+IDGM 55.22 57.45 59.85 51.41 61.34 59.41 60.61 57.90
+Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 57.85 58.42 60.40 52.50 65.22 62.05 61.42 59.69 (+1.26)
+Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
+IDGM 54.43 60.40 61.07 51.40 59.21 59.88 59.75 58.02
+Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.54 59.41 62.38 51.56 59.07 62.38 59.50 58.26 (+0.12)
+Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
+IDGM 55.49 54.92 58.36 50.31 53.15 60.35 59.21 55.97
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 56.05 56.03 59.37 52.41 53.40 62.42 60.98 57.23 (+0.45)
+Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
+IDGM 52.45 57.02 60.42 54.25 58.42 61.79 60.03 57.77
+Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 52.77 57.43 61.39 55.57 59.46 62.58 60.60 58.54 (+0.92)
+Performance of MM-IDGM We proposed MM-IDGM and conducted experiments for multi-to-single cross-domain generalization, with the results presented in Table VI.
+Both the IDGM and MM-IDGM algorithms are developed based on the domain simultaneous sampling method.
+Using the proposed Atten-Mixer fusion method as the baseline, we applied both the original IDGM and the proposed MM-IDGM.
+The results show that MM-IDGM consistently improves average accuracies across different modalities and inputs.
+However, cross-domain generalization in deception detection remains highly challenging, as some accuracies even decreased after applying IDGM.
+This may be due to the difficulty in optimizing the inner product between different domains.
+While MM-IDGM may not always perform optimally in certain cases, such as testing on the B2 dataset where dataset quality affects the results, it generally shows better performance in cross-domain generalization.
+4.6 Network Details
+Here, we present the detailed network architectures in Table VII, including encoders, the fusion module, and classifiers.
+This network architecture is utilized for experiments on multi-to-single cross-domain generalization with fusion.
+TABLE VII: Architecture details for different modalities and fusion module.
+Modality & Inputs: Visual (Face Frames)
+Input Size: [B, 3, 32, 224, 224]
+(For each frame) ResNet-18 [:-2] $\rightarrow$ [B, 32, 7, 7, 512]
+Average Pooling $\rightarrow$ [B, 32, 512]
+(For sequence) GRU Layer $\rightarrow$ [B, 32, 512]
+Fully connected layer $\rightarrow$ [B, 2]
+Modality & Inputs: Visual (AU+Gaze+Affect)
+Input Size: [B, 50, 64]
+(On d-50) Linear(50, 64) + LN + ReLU
+Linear(64, 128) + LN + ReLU
+Linear(128, 1) $\rightarrow$ [B*64, 1]
+Reshape to [B, 64, 1]
+(On d-64) Linear(64, 32) + LN + ReLU
+Linear(32, 16) + LN + ReLU
+Linear(16, 2) $\rightarrow$ [B, 2]
+Modality & Inputs: Audio (Mel Spectrogram)
+Input Size: [B, 3, 224, 224]
+ResNet-18 [:12] $\rightarrow$ [B, 7, 7, 512]
+Average Pooling $\rightarrow$ [B, 32, 512]
+Reshape to [B, 32*512]
+Linear(32*512, 12) $\rightarrow$ [B, 2]
+Atten-Mixer Fusion
+Input size: face frame [B, 32, 512],
+behavioral [B, 64, 1],
+and audio [B, 32, 512]
+Projection Layer: face frame $\rightarrow$[B, 64, 1],
+$\rightarrow$ [B, 64, 1],
+and audio $\rightarrow$ [B, 64, 1]
+Concatenation $\rightarrow$ [B, 64, 3]
+Atten-Mixer Fusion: number of patches $\rightarrow$ 64,
+channels $\rightarrow$ 3,
+and output dim $\rightarrow$ 128
+Output shape $\rightarrow$ [B, 128]
+Classifier: Linear(128, 64)+ReLU+Linear(64, 2) $\rightarrow$ [B, 2]
 4.7 Discussion
@@
 Ethical Consideration.
-Developing deception detection using AI should emphasize respecting privacy, minimizing psychological harm, preventing discrimination, promoting transparency, etc.
-Researchers should follow appropriate regulations to develop and deploy AI systems for deception detection.
-Potential misuses and negative impacts include invasion of privacy, discrimination, erosion of trust, etc.
-Mitigating these risks requires responsible practices from researchers and developers.
+Developing deception detection using AI should emphasize respecting privacy, minimizing psychological harm, preventing discrimination, etc.
+Potential misuse may cause negative impacts on society.
+If these systems are deployed without consent, they may violate individuals’ privacy rights by collecting and analyzing sensitive personal data, speech patterns, facial expressions, and body language without proper safeguards.
+Researchers should follow appropriate regulations to develop and deploy deception detection systems.
+They should collect and use data responsibly, ensuring informed consent and protecting personal information.
+The development of the models should be transparent and provide clear explanations for how decisions are made.
 5 Conclusion

### v3

@@
 1 Introduction
-Figure 1: Typical samples from different publicly deception detection datasets: Real Life Trials [1], Bag of Lies [2], MU3D [3], and Box of Lies [4].
-The samples in each row are from different datasets while those in each column are with different modalities (visual vs.
-audio) and ground truth labels (i.e., truthful vs.
+Figure 1: Some samples from six publicly available deception detection datasets: Real-Life Trial [1], Bag of Lies [2], MU3D [3], Box of Lies [4], DOLOS [5], and MDPE [6].
+Each row shows one dataset, with columns grouped by modality (visual or audio) and ground truth (truthful vs.
 deceptive).
-It can be seen that serious domain shifts (e.g., resolution/illumination/pose in visual faces and pitch/loudness/noise in audio) occur among these datasets.
+The visual samples demonstrate domain-specific differences such as resolution (e.g., high in Bag of Lies, low in Real-Life Trial), illumination, and subjects’ ethnicities.
+The audio spectrograms reveal distinct patterns: controlled datasets (MU3D, Bag of Lies, MDPE) show sharp harmonics and silent gaps, signifying stable pitch and low noise.
+Real‑Life Trial and Box of Lies display blurred harmonics and speckled noise, indicating high ambient interference.
+DOLOS shows a clear pitch with slightly elevated background noise, reflecting less controlled conditions.
 Audio-visual deception detection involves utilizing AI techniques and algorithms to automatically detect deceptive behavior in speech and facial movements [7, 8, 9, 10].
@@
 However, such an assessment can be time-consuming and require significant expertise to perform accurately.
-Recently, the development of automated deception detection systems using AI and machine learning techniques has gained significant attention as the existing methods have limitations in terms of reliability, accuracy, and scalability.
-Various multimodal datasets have been introduced, including real-life trials from court scenes [1], lab-based setups [2, 3], and game show scenarios [4].
+Recently, the development of automated deception detection systems using AI and machine learning techniques has gained significant attention as the traditional methods mentioned above have limitations in terms of reliability, accuracy, and scalability.
+Various multimodal datasets have been introduced, including real-life trials from court scenes [1], lab-based setups [2, 3, 6], and game show scenarios [4, 5].
 These datasets provide a wide variety of deceptive samples from different domains, enabling researchers to examine the effectiveness of AI models on deception detection.
-Based on these datasets, progress has been made in deception detection techniques within specific domains [7, 6, 17].
-Recent studies have utilized rich visual and audio features [18, 19, 20], such as Mel Spectrogram, emotional states, and facial action units, to enhance the performance of deception detection tasks.
+Based on these datasets, progress has been made in deception detection techniques within individual datasets where models are trained and evaluated on the same domain [9, 8, 19, 20].
+Recent studies have utilized rich visual and audio features [21, 22, 23, 19, 24, 7, 25], such as Mel Spectrogram, emotional states, and facial action units, to enhance the performance of deception detection tasks.
 However, there remains a substantial research gap that needs to be addressed.
 Specifically, fewer studies have explored the cross-domain issue, despite the presence of significant domain shifts in public deception detection datasets.
-As shown in Figure 1, domain shifts are observed in both audio and visual modalities from publicly available datasets.
-The generalizability of the models is critical for practical applications.
-Therefore, such domain shifts need to be investigated in order to develop deception detection models that can be generalized across different contexts.
-Additionally, effective methods must be proposed to alleviate the domain shift issue by fusing both audio and visual features in a meaningful way.
+As illustrated in Figure 1, the publicly available datasets used in deception detection exhibit substantial domain shifts across both visual and audio modalities.
+The differences include video resolution, illumination, subjects’ ethnicities in visuals, signal-to-noise ratios (SNRs), pitch dynamics, and background noise in audio.
+We quantified these differences by summarizing statistics for visual and audio information (details in Sec.
+4.1 and Table I).
+These domain discrepancies underline the challenge of generalizing deception detection models beyond the training domain, motivating our benchmark’s focus on cross-domain evaluation.
+Effective methods must be proposed to alleviate the domain shift issue by fusing both audio and visual features in a meaningful way.
 Addressing these issues can benefit automated deception detection systems in improving generalizability in real-world applications.
 To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets.
-Our benchmarking approach utilizes widely adopted audio and visual features, and we compare the single-to-single domain performance and multi-to-single domain generalization using different architectures.
-Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing.
+Unlike previous works that assess models only within a single dataset, our benchmark systematically evaluates cross-domain generalization using widely adopted audio and visual features through two setups: single-to-single domain generalization and multi-to-single domain generalization.
+Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented for benchmarking.
+These strategies control how batch data is sampled during training from multiple source datasets, allowing models to learn either domain-invariant or domain-specific patterns.
 To further enhance performance, we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM) to improve multi-to-single domain generalization performance and an Attention-Mixer fusion method based on MLP-Mixer [26].
 In our experiments, we compared single-to-single and multi-to-single domain generalization, incorporating three sampling strategies, one domain generalization method, and six fusion methods across two modalities.
-The benchmarking results among different features reveal that the best average fusion accuracy was 58.58% for single-to-single cross-domain generalization and 59.13% for multi-to-single cross-domain generalization.
-Moreover, for MM-IDGM, we achieved the highest average accuracy of 59.69% based on domain-simultaneous sampling.
-This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings.
+The benchmarking results among different features reveal that the best average fusion accuracy was 56.82% for single-to-single cross-domain generalization and 58.88% for multi-to-single cross-domain generalization.
+Moreover, for MM-IDGM, we achieved the highest average accuracy of 59.02% based on domain-simultaneous sampling.
+This benchmarking framework fills a critical gap in the literature and serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings.
 Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues.
 In summary, our main contributions include:
-  • Introducing a new benchmark for cross-domain audio-visual deception detection: We present a comprehensive benchmark that evaluates the generalization capacity of AI models using audio and visual modalities across multiple domains in the deception detection task.
-  • Comparing single-to-single and multi-to-single domain generalization across different architectures: We conduct extensive experiments comparing the performance of AI models in single-to-single and multi-to-single domain generalization over different architectures and modalities inputs.
-  • Testing three domain sampling strategies for multi-to-single domain generalization: We propose and evaluate three distinct domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, to assess their effectiveness and offer diverse approaches to handling domain discrepancies.
-  • Introducing Multimodal Inter-Domain Gradient Matching (MM-IDGM) for domain generalization: We propose MM-IDGM to improve multi-to-single domain generalization by aligning gradient updates between modalities, enhancing the model’s ability to transfer knowledge across different domains.
-  • Proposing the Attention-Mixer fusion method for enhanced multimodal performance: We introduce the Attention-Mixer fusion method that combines MLP-Mixer and self-attention layers to capture both intra- and inter-modal interactions better, leading to improved fusion performances in audio-visual deception detection tasks.
+  • Introducing a new benchmark for cross-domain audio-visual deception detection: We present a comprehensive benchmark designed to evaluate the model generalization using audio and visual modalities across diverse datasets with domain shifts.
+  • Comparing two evaluation protocols across different architectures: single-to-single and multi-to-single domain generalization.
+These two protocols enable comprehensive performance comparisons across varying levels of domain discrepancy.
+  • Developing and testing three domain sampling strategies for multi-to-single domain generalization: We propose and evaluate three distinct domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, to assess their effectiveness and offer diverse approaches under cross-domain settings.
+  • Introducing Multimodal Inter-Domain Gradient Matching (MM-IDGM) for domain generalization: We propose MM-IDGM to improve multi-to-single domain generalization by aligning gradient updates between audio and visual modalities, enhancing the model’s ability to transfer knowledge across different domains.
+  • Proposing the Attention-Mixer fusion method for enhanced multimodal performance: We introduce the Attention-Mixer fusion method that combines MLP-Mixer and self-attention layers to better capture both intra- and inter-modal interactions, resulting in improved multimodal integration and cross-domain generalization.
 In the rest of the paper, Sec.
@@
 Domain-by-domain: the batches are sampled from one source and then from another.
+$F_{f}$, $F_{b}$, and $F_{a}$ represent the face, behavior, and audio features, respectively.
+$L_{f}$, $L_{b}$, $L_{a}$, and $L_{g}$ denote the losses corresponding to the face, behavior, and audio classifiers, as well as the fusion layer.
 2.2 Multimodal Deception Detection
@@
 Given audio input $X_{a}$ (either Mel Spectrogram [41] or waveforms), audio features $F_{a}$ are extracted through audio encoder $\mathcal{E}_{a}$.
-The corresponding classification heads for face frames ($\mathcal{H}_{f}$), behavior features ($\mathcal{H}_{b}$), and audio features ($\mathcal{H}_{a}$) output the prediction logits $\hat{Y}_{f}$, $\hat{Y}_{b}$, and $\hat{Y}_{a}$, respectively.
-The fusion head $\mathcal{G}$ takes $F_{f}$, $F_{b}$, and $F_{a}$ as input.
+The corresponding classifiers for face frames ($\mathcal{H}_{f}$), behavior features ($\mathcal{H}_{b}$), and audio features ($\mathcal{H}_{a}$) output the prediction logits $\hat{Y}_{f}$, $\hat{Y}_{b}$, and $\hat{Y}_{a}$, respectively.
+The fusion layer $\mathcal{G}$ takes $F_{f}$, $F_{b}$, and $F_{a}$ as input.
 $\mathcal{G}$ is determined by the actual fusion method, e.g., liner layer, transformer layers, MLP, etc.
@@
 Datasets.
+TABLE I: Statistics on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), MU3D (M), DOLOS (D), and MDPE (E).
+Datasets Avg Resolution Subject Ethnicity Demographics Avg SNR (dB) Std of SNR (dB)
+Real-life Trial (R) $\sim$360×640 mix, unspecified 58, gender bias 29.30 5.13
+Bag of Lies (B1) 720×1280 100% Indian 35, gendar bias 22.63 9.31
+Box of Lies (B2) $\sim$480×640 mix, unspecified 26, (6M / 20F) 25.59 3.25
+MU3D (M) 720×1280 50% Black and 50% White 80, gender balanced 37.07 4.01
+DOLOS (D) 720×1280 mix, unspecified 213, gendar bias 30.72 4.97
+MDPE (E) 1920×1080 100% Chinese 193, gendar bias 37.17 4.28
 We benchmarked the cross-domain generalization performance based on four publicly available datasets.
-Real Life Trials [1] dataset is a popular real-world dataset collected from public court trials, which consists of 121 videos including 61 deceptive and 60 truthful video clips.
+We selected six publicly accessible, widely used multimodal deception datasets to construct our cross-domain benchmark.
+Each dataset brings unique characteristics that are beneficial for evaluating generalization.
+  • Real Life Trials [1] (R) is a popular real-world dataset collected from public court trials.
+Courtroom videos capturing genuine high-stakes deception.
+It consists of 121 videos including 61 deceptive and 60 truthful video clips.
 As it is a real-world dataset, the Real Life trial dataset has more noise on both the video and audio.
 We filtered out some corrupted videos and obtained 108 videos (54 truthful and 54 deceptive) with 58 subjects for our experiments.
-Bag of Lies [2] is a multimodal dataset collected from well-controlled lab-based scenarios, where video, audio, EEG, and gaze data are collected.
+  • Bag of Lies [2] (B1) is a multimodal dataset collected from well-controlled lab-based scenarios, where video, audio, EEG, and gaze data are collected.
 It has 35 subjects, 163 truthful and 162 deceptive video clips.
 The backgrounds for the videos are relatively clean and it is less noisy.
-MU3D [3] has 320 video clips and 80 subjects that cover different races and genders.
+  • MU3D [3] (M) has 320 video clips and 80 subjects that cover different races and genders.
 It is also a lab-based dataset that uses the personal description paradigm to stimuli real-world cases.
 Each participant tells a positive truth, a positive lie, a negative truth, and a negative lie.
-Box of Lies [4] is a deception dataset collected from an online gameshow, which has 25 videos and 26 participants (6 male and 20 female).
+  • Box of Lies [4] (B2) is a deception dataset collected from an online gameshow, which has 25 videos and 26 participants (6 male and 20 female).
 The full video set contains 29 truthful and 36 deceptive rounds of games.
@@
 After preprocessing, 101 video clips were extracted for testing.
+  • DOLOS [5] (D) is the largest game show deception detection dataset, featuring rich deceptive conversations.
+It consists of 1,675 clips from 213 subjects, including 899 deceptive and 776 truthful samples, with fine-grained audio-visual annotations.
+The dataset is notable for its higher quality compared to existing alternatives.
+  • MDPE [6] (E) contains over 104 hours of deception content with 193 subjects.
+It supports not only deception detection but also downstream tasks such as personality recognition, emotion analysis.
+It is a lab-based dataset with monetary incentives to encourage naturalistic deceptive behaviors.
+Each video has a 5-grade truthful to deceptive labels from the subjects.
+This enables analysis of subject-level variation in deception behavior.
+Taken together, these six datasets present diverse domains in terms of subject demographics, recording conditions, modalities, and data quality.
+This diversity is essential to rigorously evaluate cross-domain generalization for multimodal deception detection.
 Some of the typical samples from these datasets are shown in Fig.
-1.
+1 and more statistics are provided in Table I.
 Evaluation Metrics.
 In this work, we followed the widely adopted metric, binary classification accuracy (%), for experimental evaluation.
-The deceptive clips were labeled as 1 and truthful clips were labeled as 0.
+The deceptive clips were labeled as 1 and the truthful clips were labeled as 0.
+TABLE II: The results of intra-domain testing accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), MU3D (M), DOLOS (D), and MDPE (E).
+Modality & Inputs Method R to R B1 to B1 M to M D to D E to E
+Visual (Gaze) MLP 59.96 53.45 52.03 54.95 51.89
+Visual (AU+Gaze+Affect) MLP 62.87 55.84 55.23 57.18 55.38
+Visual (Facial Landmark) LSTM 58.32 50.07 51.24 52.03 51.26
+Visual (Face frames) ResNet18 70.10 57.78 55.15 65.23 56.57
+Visual (Face frames) ResNet18+KNN 65.28 53.68 52.04 61.36 54.92
+Visual (Face frames) ResNet18+SVM 66.52 54.21 54.36 62.17 56.08
+Visual (Face frames) FFCSN 67.33 56.82 56.93 63.26 57.80
+Audio (Acoustic + Prosodic) MLP 66.82 50.49 55.31 58.96 50.47
+Audio (Mel spectrogram) ResNet18 75.78 55.69 56.45 60.06 52.99
+Audio (Mel spectrogram) ResNet18+KNN 73.24 56.38 57.22 59.07 52.30
+Audio (Mel spectrogram) ResNet18+SVM 73.25 56.88 58.04 60.11 52.69
+concat 59.09 54.46 56.30 70.40 55.78
+Visual (Face frames) KNN 60.22 54.83 70.21 69.50 56.22
++ Visual (AU+Gaze+Affect) SVM 61.03 55.21 69.96 70.05 55.39
+Atten-Mixer (Ours) 63.05 55.45 57.59 71.55 56.18
+concat 68.75 54.74 56.19 66.95 53.39
+Visual (Face frames) KNN 67.33 55.32 56.62 65.98 53.16
++ Audio (Mel spectrogram) SVM 68.45 56.08 56.36 67.45 54.03
+Atten-Mixer (Ours) 71.00 56.44 57.14 68.10 54.58
+concat 68.88 56.25 53.27 68.97 51.39
+Visual (AU+Gaze+Affect) KNN 67.92 54.23 50.32 65.23 50.46
++ Audio (Mel spectrogram) SVM 68.03 54.50 51.06 64.92 51.35
+Atten-Mixer (Ours) 69.30 56.78 55.23 70.98 52.99
+concat 69.33 58.43 55.24 72.70 54.98
+Visual (Face frames) KNN 68.55 54.25 52.28 68.48 53.19
++ Visual (AU+Gaze+Affect) SVM 68.22 54.39 52.47 68.65 54.24
++ Audio (Mel spectrogram) Atten-Mixer(Ours) 70.52 59.41 56.88 73.85 58.17
+TABLE III: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), MU3D (M), DOLOS (D), and MDPE (E).
+Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 D to R E to R E to B1 E to B2
+Visual (AU) LSTM 48.11 - - 61.21 - - - - - - - - -
+Visual (Gaze) MLP 47.22 59.98 54.15 57.23 55.42 49.06 45.38 50.35 52.22 53.08 50.16 51.45 53.07
+Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79 51.29 52.33 54.68
+Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58 51.36 52.07 54.88
+Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41 51.89 52.77 55.45
+Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52 51.02 53.18 55.92
+Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80 52.03 55.06 56.32
+Visual (Face frames) ResNet18+KNN 53.13 63.15 51.72 56.38 59.25 51.06 47.02 53.22 56.11 55.36 51.96 55.88 56.15
+Visual (Face frames) ResNet18+SVM 53.66 63.28 52.13 56.04 60.24 52.15 47.88 53.26 55.67 56.36 52.14 55.39 56.17
+Visual (Face frames) FFCSN 54.32 62.88 51.06 59.36 53.11 48.25 52.65 52.17 54.36 56.85 53.36 54.88 56.37
+Audio (Acoustic + Prosodic) MLP 43.06 52.18 50.36 49.44 59.56 51.46 50.17 50.45 52.04 50.38 46.98 49.77 50.03
+Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43 50.56 52.04 53.68
+Audio (Mel spectrogram) ResNet18+KNN 46.88 54.02 51.39 50.69 62.77 49.36 51.09 51.53 52.72 53.17 50.48 51.72 53.23
+Audio (Mel spectrogram) ResNet18+SVM 49.98 54.02 51.36 50.45 62.17 50.33 52.04 52.48 53.66 54.25 51.26 52.33 54.05
+Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55 52.77 53.49 55.68
+Modality & Inputs Method R to D R to E B1 to D B1 to E M to D M to E D to B1 D to B2 D to M D to E E to M E to D Avg
+Visual (AU) LSTM 49.92 - - - - - - - - - - - 53.08
+Visual (Gaze) MLP 49.98 57.46 50.21 50.55 52.02 49.87 55.36 50.11 53.29 52.41 49.32 50.21 51.98
+Visual (AU+Gaze) MLP 51.25 52.33 51.64 53.33 52.47 50.66 50.71 51.37 53.68 55.22 50.04 51.42 53.21
+Visual (Affect) MLP 51.32 53.35 52.56 53.47 50.98 51.42 51.34 51.08 54.37 54.56 51.21 51.49 53.25
+Visual (AU+Gaze+Affect) MLP 52.38 54.24 53.19 53.98 51.32 52.26 51.20 50.25 55.08 55.02 51.26 51.49 53.84
+Visual (Face frames) ResNet18 52.75 51.28 52.18 50.65 52.32 52.36 50.89 51.33 55.42 54.62 51.25 52.08 53.30
+Visual (Face frames) ResNet18+GRU 52.54 52.36 53.88 50.87 52.68 53.94 51.34 52.36 55.69 55.22 52.36 51.44 53.84
+Visual (Face frames) ResNet18+KNN 53.37 62.32 51.26 56.37 59.04 50.77 48.21 51.48 54.39 55.08 51.49 52.33 54.26
+Visual (Face frames) ResNet18+SVM 54.15 61.28 50.36 55.49 59.68 50.28 49.47 50.28 53.66 54.88 52.74 54.29 54.44
+Visual (Face frames) FFCSN 53.87 61.46 51.42 56.23 60.13 50.35 50.31 53.19 53.62 54.07 53.28 55.03 54.50
+Audio (Acoustic + Prosodic) MLP 45.58 51.04 51.26 50.33 60.82 49.92 51.83 50.22 53.16 50.59 49.25 50.21 50.80
+Audio (Mel spectrogram) ResNet18 50.01 52.11 50.59 50.48 60.37 51.29 53.08 53.86 54.36 53.68 50.26 51.42 52.88
+Audio (Mel spectrogram) ResNet18+KNN 48.82 49.98 50.21 50.28 60.08 50.49 54.11 52.38 54.59 54.32 50.38 51.27 52.24
+Audio (Mel spectrogram) ResNet18+SVM 49.05 50.31 49.89 60.49 56.75 52.98 55.32 53.98 55.84 54.28 51.26 52.38 53.24
+Audio (Waveform) Wave2Vec 50.21 49.51 49.94 50.43 55.83 50.82 54.26 53.65 54.38 53.49 51.21 50.65 52.87
 4.2 Implementation Details
@@
 Affect features were extracted by Emonet [40], where the feature included 5-class emotions, arousal, and valence.
-For audio features, Mel Spectrograms were extracted by OpenSmile toolkit [38].
+For audio features, Mel Spectrograms and acoustic features were extracted by the OpenSmile toolkit [41].
 Raw audio waveforms were also used in our experiments.
@@
 To evaluate the models’ cross-domain generalization capacity and alleviate domain information leakage, all the preprocessed data including original training and test data from each dataset was used for either training or testing.
-Note that the Box of Lies dataset was only used for testing as many samples were filtered out due to their unsatisfactory quality.
+Note that the Box of Lies dataset was only used for testing, as many samples were filtered out due to their unsatisfactory quality.
+Because of shifting camera angles and background audience noise, the speaker’s face and speech may not be captured clearly in the video.
 The experiments were conducted on the single-to-single domain (e.g., R to B1 stands for training on Real-life Trial (R) and testing on Bag of Lies (B1)) and multi-to-single domain (e.g., R&M to B2 stands for training on Real-life Trial (R) and MU3D (M) and testing on Box of Lies (B2)).
 Model Selection.
+Different audio and visual features are used for benchmarking, including gaze, AUs, affect, facial landmark (2D, 68 points, using dlib package), acoustic and prosodic features (88-dimensional features extracted from the OpenSmile toolkit), Mel spectrograms, and audio waveforms.
+We also benchmark and re-implement the methods that are widely adopted by domain-specific deception detection task [20], such as ResNet18, MLP, SVM (Linear) [19, 25], KNN [24], FFCSN [7], LSTM [19], etc.
 Models for audio and visual modalities were selected to fit the data volume.
-For face frames, we adopted ResNet18 [35] and Gate Recurrent Unit (GRU) [46] models for facial feature extraction and temporal modeling, respectively.
-Two-layer multilayer perception (MLP) [47] models were used for AUs, gaze, and affect feature representation.
+For face frames, we adopted ResNet18 [38], and Gate Recurrent Unit (GRU) [49], SVM [19, 25], and KNN [24] models for facial feature extraction and temporal modeling, respectively.
+Two-layer multilayer perception (MLP) [50] or LSTM [19] models were used for AUs, gaze, and affect feature representation.
 Specifically, the first MLP Layer initially operates on the dimension that represents the number of features, with sizes following the sequence: “input dimension $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 1.
@@
 The MLP networks have identical structures, except for their first linear layers, as the input dimensions differ.
-For the audio-based Mel spectrogram, we used the ResNet18 [35] model for time-frequency feature representation.
+For acoustic features, MLP is used.
+For the audio-based Mel spectrogram, we used the ResNet18 [38], KNN, and SVM models for time-frequency feature representation and classification.
 For audio waveforms, the Wave2Vec [51] model was applied for audio feature extraction.
@@
 The models were trained with batchsize 16 for a maximum of 30 epochs.
-TABLE I: The results of intra-domain testing accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), and MU3D (M).
-Modality & Inputs Method R to R B1 to B1 M to M
-Visual (Face frames) ResNet18 70.10 57.78 55.15
-Visual (AU+Gaze+Affect) MLP 62.87 55.84 55.23
-Audio (Mel spectrogram) ResNet18 75.78 55.69 56.45
-Visual (Face frames) concat 59.09 54.46 56.30
-+ Visual (AU+Gaze+Affect) Atten-Mixer (Ours) 63.05 55.45 57.59
-Visual (Face frames) concat 68.75 54.74 56.19
-+ Audio (Mel spectrogram) Atten-Mixer (Ours) 71.00 56.44 57.14
-Visual (AU+Gaze+Affect) concat 68.88 56.25 53.27
-+ Audio (Mel spectrogram) Atten-Mixer (Ours) 69.30 56.78 55.23
-Visual (Face frames) concat 69.33 58.43 55.24
-+ Visual (AU+Gaze+Affect) Atten-Mixer(Ours) 70.52 59.41 56.88
-+ Audio (Mel spectrogram)
-TABLE II: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
-Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
-Visual (AU) LSTM 48.11 - - 61.21 - - - - - -
-Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52
-Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80
-Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79
-Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58
-Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41
-Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43
-Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55
-TABLE III: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
-Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
+TABLE IV: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
 Domain-Simultaneous
-Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 50.51
-Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 52.53
-Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 51.82
-Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 55.04
-Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.4 63.37 55.82
-Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.5 51.12
-Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 51.78
+Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
+Visual (Face frames) ResNet18 53.85 49.50 49.50 50.94 44.86 60.40 44.55 52.43 50.88 54.82 49.92
+Visual (Face frames) ResNet18+GRU 52.62 54.46 51.49 51.88 53.27 59.41 44.55 53.06 51.88 53.98 50.33
+Visual (Face frames) ResNet18+SVM 52.98 54.87 52.08 52.66 54.12 59.39 45.38 53.56 52.39 54.08 51.24
+Visual (AU+Gaze) MLP 53.54 47.52 48.51 50.94 52.34 53.47 56.44 50.48 51.28 53.35 50.22
+Visual (Affect) MLP 50.15 54.46 55.45 52.19 53.27 57.43 62.38 52.08 52.33 53.24 51.02
+Visual (AU+Gaze+Affect) MLP 50.46 52.48 61.39 51.25 51.4 60.40 63.37 52.92 53.45 53.99 52.27
+Audio (Mel spectrogram) ResNet18 48.92 45.54 53.47 53.12 43.93 62.38 50.50 51.06 50.88 52.25 51.43
+Audio (Mel spectrogram) ResNet18+SVM 49.88 47.99 52.64 52.96 46.98 59.88 51.28 52.04 51.36 53.82 52.77
+Audio (Waveform) Wave2Vec 52.92 55.45 44.55 51.25 69.16 42.57 46.53 53.26 52.18 53.69 53.06
+Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Visual (Face frames) ResNet18 50.21 52.39 51.09 51.39 49.55 50.03 51.88 50.58 51.84 50.79 51.02
+Visual (Face frames) ResNet18+GRU 51.25 52.98 52.92 52.24 50.26 51.26 52.26 51.23 52.06 51.26 52.13
+Visual (Face frames) ResNet18+SVM 52.56 53.26 52.36 53.62 51.22 51.48 53.18 51.46 52.47 51.63 52.67
+Visual (AU+Gaze) MLP 50.48 52.48 52.02 51.98 50.23 51.06 51.66 50.23 51.02 50.16 51.40
+Visual (Affect) MLP 51.35 53.45 52.88 53.26 52.45 52.48 52.87 51.45 52.38 51.26 53.23
+Visual (AU+Gaze+Affect) MLP 52.48 54.28 54.33 54.18 52.35 53.18 61.46 53.29 54.39 53.28 54.60
+Audio (Mel spectrogram) ResNet18 51.33 51.47 52.02 51.05 50.25 50.49 51.06 51.38 50.29 51.33 51.15
+Audio (Mel spectrogram) ResNet18+SVM 52.55 52.39 52.78 52.49 51.62 51.63 52.68 52.58 51.64 52.68 52.13
+Audio (Waveform) Wave2Vec 53.09 53.69 53.85 52.63 52.04 51.38 52.04 52.44 51.83 52.82 52.40
 Domain-Alternating
-Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 53.50
-Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 56.25
-Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 54.25
-Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 55.05
-Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 55.66
-Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 53.17
-Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.57
+Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
+Visual (Face frames) ResNet18 50.15 45.54 56.44 51.56 50.47 54.46 65.85 52.28 51.34 54.65 50.32
+Visual (Face frames) ResNet18+GRU 55.38 52.48 60.40 50.00 50.47 60.40 64.62 53.66 52.26 54.03 51.29
+Visual (Face frames) ResNet18+SVM 56.21 52.48 61.23 51.05 51.28 59.39 64.28 53.54 53.65 54.88 52.43
+Visual (AU+Gaze) MLP 55.45 47.52 53.47 51.25 54.21 57.43 60.40 51.29 52.26 53.46 50.33
+Visual (Affect) MLP 51.08 56.44 61.39 52.19 52.34 58.42 53.47 52.45 52.56 53.83 50.98
+Visual (AU+Gaze+Affect) MLP 51.38 58.42 63.37 50.31 53.27 52.48 60.40 53.76 53.58 54.72 51.26
+Audio (Mel spectrogram) ResNet18 50.15 60.40 53.47 50.31 58.88 51.49 47.52 52.02 51.63 54.22 50.36
+Audio (Mel spectrogram) ResNet18+SVM 51.53 59.62 53.02 51.26 60.38 52.56 47.33 51.76 51.38 53.55 51.82
+Audio (Waveform) Wave2Vec 52.92 55.45 44.55 50.62 64.49 58.42 48.51 53.26 52.35 53.29 52.47
+Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Visual (Face frames) ResNet18 52.34 52.88 51.38 51.56 51.42 50.48 51.25 51.25 51.66 50.39 52.27
+Visual (Face frames) ResNet18+GRU 52.11 54.04 53.25 53.43 51.04 52.22 53.26 52.18 52.36 54.32 53.96
+Visual (Face frames) ResNet18+SVM 52.66 54.65 52.89 53.75 51.02 51.46 52.35 52.21 53.07 53.32 54.18
+Visual (AU+Gaze) MLP 51.25 52.02 53.46 52.33 51.75 51.21 52.15 50.52 51.32 50.88 52.75
+Visual (Affect) MLP 51.62 52.72 53.86 54.33 51.02 52.38 51.77 52.63 51.56 51.74 53.28
+Visual (AU+Gaze+Affect) MLP 52.02 54.32 52.64 54.31 51.56 53.22 52.45 53.25 52.36 52.41 53.88
+Audio (Mel spectrogram) ResNet18 50.21 51.62 51.32 52.46 50.28 50.11 51.24 51.36 50.86 52.46 52.02
+Audio (Mel spectrogram) ResNet18+SVM 52.32 51.85 52.03 51.24 50.38 51.26 52.37 52.44 51.69 51.76 52.45
+Audio (Waveform) Wave2Vec 52.94 53.26 52.68 53.06 52.66 52.46 52.10 52.21 51.62 52.47 52.94
 Domain-by-Domain
-Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 52.68
-Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.46
-Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 52.68
-Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 55.05
-Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 56.67
-Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 54.24
-Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 54.90
-TABLE IV: The fusion results of single-to-single cross-domain generalization accuracy (%).
-Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 Avg
-Score-level Average 53.23 41.58 51.88 64.49 62.38 50.62 57.94 53.23 62.38 55.30
-Concat 51.08 55.45 51.88 54.21 58.42 51.25 57.01 50.15 61.39 54.54
-SE-Concat 53.85 60.40 51.25 55.14 58.42 50.62 56.07 51.38 65.35 55.83
-Cross-Atten 55.38 61.39 52.19 51.40 60.40 51.25 55.14 56.31 60.40 55.98
-MLP-Mixer 55.08 48.51 53.44 56.07 58.42 53.75 55.14 59.08 59.41 55.43
-Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.92 59.41 57.94 63.37 53.75 53.75 60.75 56.00 61.39 58.14
-Score-level Average 53.23 49.50 51.88 50.47 59.41 53.75 65.42 56.31 49.50 54.39
-Concat 50.77 53.47 50.47 62.38 51.56 51.88 54.21 52.62 58.42 53.98
-SE-Concat 50.15 44.55 51.40 61.39 53.12 52.50 65.42 56.31 66.34 55.69
-Cross-Atten 54.46 51.49 55.14 58.42 51.25 52.19 63.55 56.62 66.34 55.95
-MLP-Mixer 52.31 55.45 57.94 63.37 53.12 51.25 64.49 57.85 62.38 57.57
-Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.54 55.45 56.07 61.39 50.94 53.12 67.29 57.85 64.36 58.22
-Score-level Average 49.85 58.42 54.06 45.79 60.40 50.62 49.53 56.00 63.37 54.23
-Concat 49.54 53.47 47.66 61.39 53.44 51.88 57.01 50.15 58.42 53.66
-SE-Concat 50.15 48.51 55.14 60.40 53.12 50.00 57.01 49.85 63.37 54.17
-Cross-Atten 53.23 44.55 57.94 55.45 54.69 54.06 63.55 51.69 64.36 55.50
-MLP-Mixer 49.54 57.43 50.47 63.37 54.06 53.12 59.81 55.08 69.31 56.91
-Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.52 54.46 57.94 61.39 53.12 51.56 69.16 58.15 64.36 58.18
-Score-level Average 52.00 58.42 51.88 53.27 58.42 50.94 57.01 53.54 61.39 55.20
-Concat 53.23 58.42 55.14 61.39 52.50 52.19 56.07 54.15 60.40 55.94
-SE-Concat 51.38 58.42 51.40 62.38 52.81 50.94 59.81 54.77 52.48 54.93
-Cross-Atten 51.08 48.51 55.14 60.40 53.12 53.44 60.75 56.31 60.40 55.46
-MLP-Mixer 55.69 46.53 44.86 63.37 51.56 50.94 64.49 56.00 60.40 54.87
-Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 60.40 57.01 64.36 53.44 51.25 67.29 56.00 62.38 58.58
+Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
+Visual (Face frames) ResNet18 52.00 53.47 56.44 50.00 59.81 41.58 55.45 54.28 53.22 53.27 52.88
+Visual (Face frames) ResNet18+GRU 54.46 41.58 66.34 50.62 51.40 56.44 60.40 54.69 52.59 53.64 53.48
+Visual (Face frames) ResNet18+SVM 55.25 51.36 62.17 52.33 51.28 52.54 53.31 52.42 53.88 54.21 52.31
+Visual (AU+Gaze) MLP 51.08 43.56 55.45 53.75 57.01 53.47 54.46 51.52 52.43 53.66 51.85
+Visual (Affect) MLP 55.69 57.43 57.43 51.56 52.34 49.50 61.39 53.11 53.00 54.28 52.14
+Visual (AU+Gaze+Affect) MLP 50.15 56.44 58.42 50.00 57.94 60.40 63.37 54.31 53.26 53.65 53.15
+Audio (Mel spectrogram) ResNet18 52.31 50.50 58.42 49.38 53.27 56.44 59.41 51.06 52.31 51.45 52.06
+Audio (Mel spectrogram) ResNet18+SVM 51.36 50.67 52.31 50.42 52.57 56.38 58.01 51.88 52.04 52.35 52.38
+Audio (Waveform) Wave2Vec 56.00 47.52 44.55 53.12 67.29 57.43 58.42 52.06 52.46 53.06 53.22
+Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Visual (Face frames) ResNet18 52.14 53.02 52.69 51.49 56.33 52.46 52.72 52.04 51.48 51.31 52.77
+Visual (Face frames) ResNet18+GRU 52.68 53.06 55.38 55.64 55.58 56.03 55.42 52.35 53.92 52.65 54.21
+Visual (Face frames) ResNet18+SVM 53.03 55.25 55.06 54.23 56.63 58.02 55.37 53.24 52.85 54.31 54.24
+Visual (AU+Gaze) MLP 51.94 52.31 52.58 54.62 55.31 53.27 52.48 51.56 52.47 51.78 52.69
+Visual (Affect) MLP 52.15 54.36 55.40 53.19 52.27 55.43 52.52 51.23 50.82 51.25 53.64
+Visual (AU+Gaze+Affect) MLP 52.39 53.48 54.28 53.62 52.25 53.56 53.06 52.18 51.59 52.46 54.28
+Audio (Mel spectrogram) ResNet18 54.24 55.23 52.63 55.82 53.22 55.21 52.89 51.28 53.97 56.22 53.68
+Audio (Mel spectrogram) ResNet18+SVM 52.26 54.98 53.48 55.28 52.01 53.83 53.09 52.35 55.22 51.09 53.05
+Audio (Waveform) Wave2Vec 54.99 55.93 53.29 55.83 54.25 53.29 53.26 53.25 55.42 56.99 54.36
 4.3 Cross-domain Testing with Unimodal Features
@@
 Intra-Domain Testing.
-Before performing cross-domain testing, we show the results of intra-domain testing on Real-life Trial (R), Bag of Lies (B1), and MU3D (M) in Table I.
+Before performing cross-domain testing, we show the results of intra-domain testing on Real-life Trial (R), Bag of Lies (B1), MU3D (M), DOLOS (D), and MDPE (E) in Table II.
 It is important to note that we did not use the Box of Lies (B2) dataset for training due to concerns about its quality.
-Specifically, we performed 5-fold intra-domain testing and reported the average accuracies.
-ResNet18 and MLP were used as baseline models, as they are widely used as encoders in audio-visual deception detection.
-As expected, the performance of intra-domain testing varies across datasets, reflecting the differences in their levels of difficulty.
+Specifically, we performed a subject-independent five-fold cross-testing and reported the average accuracies.
+Among unimodal inputs, mel spectrograms with ResNet18 achieve the highest accuracy (75.78% on R), highlighting the strength of audio cues, particularly in structured settings like courtroom videos.
+Visual face frames also perform well with higher average results compared with audio features.
+However, facial landmarks are less reliable and show consistently low accuracy, which may due to the sparse geometric information and its sensitivity to pose, occlusion, and lighting; thus, we exclude them from the following cross-domain tasks.
+Multimodal fusion consistently improves performance, with the proposed Atten-Mixer achieving the best results across several datasets (e.g., 73.85% on D, 58.17% on E).
+Performance variation across datasets suggests that data quality, recording conditions, and content type significantly impact model effectiveness.
+Overall, the intra-testing results indicate the benefit of combining rich audio and visual features with advanced fusion methods for improved deception detection performances.
 Single-to-Single Domain.
-Specifically, the models were trained on one dataset from one domain and tested on the other dataset from another domain.
-The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
-As shown in Tabel II, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect.
-For audio modality, Mel spectrogram and waveform were extracted.
-We also applied several different backbone networks as audio and visual encoders.
-We can observe that R and B1 datasets generalized the best on B2, and M generalized the best on R.
-Note that the B2 dataset was not adopted as a source domain dataset as the original dataset had too much noise and we cleaned it only for testing.
-On average, we can observe that the best result was achieved by using visual (AU+gaze+affect) features.
+Specifically, the models were trained on one dataset (one domain) and tested on the other dataset (another domain).
+As shown in Tabel III, overall, visual-based models exhibit strong generalization compared to audio-only models, particularly those leveraging rich visual descriptors (AU, gaze, affect) or deep networks like ResNet18+KNN and FFCSN.
+Notably, the best average performance across all transfer scenarios is achieved by FFCSN (54.50%), closely followed by visual-only model ResNet18+SVM (54.44%) and ResNet18+KNN (54.26%), highlighting the effectiveness of visual temporal modeling.
+Although audio contributes valuable cues, it struggles more with domain shifts compared to visual modalities.
+Additionally, handcrafted prosodic/acoustic features perform worse overall (e.g., MLP: 50.80% avg), compared with other types of audio features.
 Multi-to-Single Domain.
-Here we fully evaluate the performance of multi-to-single cross-domain generalization using unimodal features.
-We conducted experiments for different domain sampling strategies.
-As shown in Table III, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%).
-For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%).
-The results showed that the best generalization performances were obtained when transferring to the real-world dataset and the gameshow dataset by training on the lab-based datasets.
-This was because the lab-based datasets (Bag of Lies and MU3D) are relatively clean compared to the real-world dataset (Real-life Trial) and the gameshow dataset (Box of Lies).
-However, in the opposite case, the generalization performance degraded, for example, R&M to B1 and R&B1 to M.
-Different domain sampling strategies reached their best average performance on different input features and backbone networks.
-To be specific, for both domain-simultaneous and domain-by-domain strategies, models trained on visual (AU+gaze+affect) features reached their highest accuracies, which were 55.82% and 56.67%, respectively.
-Using the domain-alternating strategy, the best accuracy of 56.25% was achieved by training on visual (face frames) features.
-We can observe that models trained on visual modalities outperformed those trained on audio modalities across all the generalization strategies.
-This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
-Discussion: (1) Wave2Vec outperforms MFCC in audio feature extraction due to its ability to capture richer, more meaningful representations, as demonstrated across various sampling methods.
-Additionally, larger and more diverse datasets like B1 and M enhance generalization, leading to improved performance when models trained on these datasets are tested on smaller or less challenging datasets like R.
-The R dataset, composed of courtroom trial videos, tends to yield better results, while B2, a game show dataset, often produces unsatisfactory outcomes due to its lower video quality and data imbalance.
-Overall, visual features perform better than audio features and models using Wave2Vec features achieve better results compared to those trained on MFCC.
-(2) While domain-simultaneous sampling is typically expected to enhance generalization by promoting domain-invariant learning, our average results show that domain-by-domain sampling slightly outperforms it in these audio-visual deception detection tasks.
-This may be due to its ability to capture domain-specific variations, which are beneficial in this context.
-However, the effectiveness of each sampling strategy depends on the features being used; for example, Wave2Vec features perform better with domain-simultaneous sampling, while MFCC features benefit more from domain-by-domain sampling.
-This suggests that the optimal sampling strategy may vary depending on the feature set’s predictive power and the specific dataset characteristics.
-(3) Visual features, particularly behavioral features like AUs, gaze, and affective cues, outperform audio features in most of our experiments.
-Behavioral features are more predictive in the majority of testing pairs, as seen in Table II, and achieved the best average performance across sampling strategies in Table III.
-Their rich information content, capturing subtle signs of deception, may explain their superior performance over the face or audio features.
-Due to the lower dimension compared with face and audio features, behavioral features make them more effective when training on small datasets for detecting deception.
+Table IV presents the domain generalization results under three sampling strategies: Domain-Simultaneous, Domain-Alternating, and Domain-by-Domain.
+Each strategy evaluates 21 domain generalization tasks using a consistent set of nine input features and model configurations.
+The final row in each strategy block reports the average accuracy across all 21 tasks, serving as an overall performance indicator for that sampling strategy.
+It is worth noting that these 21 tasks do not represent the full combination of the six datasets, but they cover most scenarios and are sufficiently comprehensive for evaluation.
+Our results indicate that the sampling strategy, feature selection, and dataset composition all play critical roles in building generalizable deception detection systems.
+We offer several key observations and discussions below:
+(1) Among the three strategies, Domain-by-Domain achieves the best overall average accuracy (average across all the results) of 53.66%, followed by Domain-Alternating (53.06%) and Domain-Simultaneous (52.30%).
+This suggests that training on individual domains sequentially allows models to better abstract transferable patterns, likely due to reduced complexities between domains during training.
+In contrast, domain-alternating and domain-simultaneous keep switching between source domains, appearing to induce higher variability and lower overall accuracies.
+(2) Across all strategies, models trained on visual features generally outperform those trained on audio features.
+This is shown in the average results from each block.
+For domain-simultaneous training, the best performance (54.60%) is achieved using visual behavioral features (AU+Gaze+Affect) with an MLP classifier.
+For domain-alternating, the best average performance is achieved by visual (face frames) trained with ResNet18+SVM.
+For domain-by-domain, although the best performance is achieved by audio (Waveform), its visual performances outperform audio ones, highlighting the robustness of visual cues over audio modalities.
+(3) Regarding different features, overall, visual and behavioral features perform better than audio features.
+However, the performance using different strategies varies.
+For domain-simultaneous, we can observe that behavioral features (AUs, gaze, and affective cues) are more predictive than face frames or audio features.
+Their lower dimensionality and interpretability allow effective learning even on limited training data.
+For domain-alternating, face frames is better.
+For domain-by-domain, audio features achieve the best performance.
+Wave2Vec generally outperforms handcrafted features like Mel spectrograms or acoustic-prosodic embeddings, highlighting the importance of using pretrained self-supervised representations.
+(4) Regarding dataset effects, one observation is that training on clean, lab-collected datasets (such as Bag of Lies and MU3D) tends to generalize better to real-world or noisier datasets (like Real-life Trial or Box of Lies).
+For example, under domain-simultaneous sampling, training on B1&M and testing on R yields the best result of 69.16%.
+This performance is consistent with the domain-by-domain strategy, where B1&M to R reaches 67.29%.
+Conversely, when training noisy or lower-quality datasets (e.g., R, E), performance often degrades, suggesting that domain complexity and data quality significantly affect generalization.
+Figure 3: Performance comparisons of Domain-simultaneous training w/ and w/o Gradient Reversal Layer (GRL).
 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
-Following the implementation by Ganin et al.
-[49], we compared the multi-to-single domain generalization accuracies w/w.o GRL.
+Gradient Reversal Layer (GRL) [52] has become a foundational component in domain adaptation and domain generalization due to its ability to encourage domain-invariant feature learning.
+Its effectiveness and versatility have been demonstrated across diverse fields such as face anti-spoofing [42], sleep stage classification [53], multi-view 3D object detection [54], etc.
+Following the implementation by Ganin and Lempitsky [52], we compared the multi-to-single domain generalization accuracies w/w.o GRL.
 GRL was proposed to mitigate the domain shift issue by manipulating the training gradients.
@@
 Generally, ResNet18 trained with GRL performed better than MLP for visual modality, and Wave2Vec trained with GRL boosted the performance and surpassed the model trained on the Mel spectrogram.
-Figure 3: Performance comparisons of Domain-simultaneous training w/ and w/o Gradient Reversal Layer (GRL).
-TABLE V: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
-Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
+TABLE V: Fusion results of single-to-single cross-domain generalization accuracy (%).
+Modality & Inputs Fusion Postion Fusion Method R to B1 R to B2 R to M R to D R to E B1 to R B1 to B2 B1 to M B1 to D B1 to E
+Score-level Average 53.23 41.58 51.88 53.85 50.16 64.49 62.38 50.62 52.29 50.97
+Concat 51.08 55.45 51.88 52.90 50.49 54.21 58.42 51.25 51.94 50.08
+SE-Concat 53.85 60.40 51.25 53.07 51.29 55.14 58.42 50.62 55.32 50.81
+Cross-Atten 55.38 61.39 52.19 54.11 52.35 51.40 60.40 51.25 56.09 51.05
+MLP-Mixer 55.08 48.51 53.44 54.02 52.18 56.07 58.42 53.75 57.22 51.29
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.92 59.41 57.94 55.32 53.22 63.37 53.75 53.75 57.39 51.38
+Score-level Average 53.23 49.50 51.88 53.69 50.81 50.47 59.41 53.75 52.90 50.49
+Concat 50.77 53.47 50.47 52.64 50.24 62.38 51.56 51.88 51.69 49.35
+SE-Concat 50.15 44.55 51.40 51.86 50.16 61.39 53.12 52.50 54.02 50.24
+Cross-Atten 54.46 51.49 55.14 53.93 51.21 58.42 51.25 52.19 54.80 50.49
+MLP-Mixer 52.31 55.45 57.94 54.36 51.54 63.37 53.12 51.25 55.66 51.21
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.54 55.45 56.07 55.06 51.78 61.39 50.94 53.12 57.13 52.35
+Score-level Average 49.85 58.42 54.06 52.16 50.08 45.79 60.40 50.62 51.08 50.57
+Concat 49.54 53.47 47.66 50.91 50.57 61.39 53.44 51.88 50.39 50.49
+SE-Concat 50.15 48.51 55.14 50.99 50.65 60.40 53.12 50.00 52.93 50.24
+Cross-Atten 53.23 44.55 57.94 51.25 50.89 55.45 54.69 54.06 55.14 51.38
+MLP-Mixer 49.54 57.43 50.47 53.33 51.38 63.37 54.06 53.12 56.18 51.70
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.52 54.46 57.94 54.97 51.13 61.39 53.12 51.56 57.48 52.43
+Score-level Average 52.00 58.42 51.88 52.43 50.97 53.27 58.42 50.94 54.11 50.16
+Concat 53.23 58.42 55.14 53.93 50.32 61.39 52.50 52.19 55.23 50.32
+SE-Concat 51.38 58.42 51.40 53.59 50.24 62.38 52.81 50.94 56.53 51.62
+Cross-Atten 51.08 48.51 55.14 54.11 51.38 60.40 53.12 53.44 57.22 50.08
+MLP-Mixer 55.69 46.53 44.86 54.97 51.05 63.37 51.56 50.94 57.04 51.21
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 60.40 57.01 56.35 51.86 64.36 53.44 51.25 57.56 51.62
+Modality & Inputs Fusion Postion Fusion Method M to R M to B1 M to B2 M to D M to E D to R D to B1 D to B2 D to M D to E
+Score-level Average 57.94 53.23 62.38 53.41 50.32 52.31 51.69 52.28 50.46 50.32
+Concat 57.01 50.15 61.39 53.50 51.21 53.84 50.46 54.08 50.40 50.16
+SE-Concat 56.07 51.38 65.35 54.19 50.57 54.42 51.38 54.48 50.37 50.77
+Cross-Atten 55.14 56.31 60.40 54.39 51.13 55.23 52.62 57.43 50.86 50.97
+MLP-Mixer 55.14 59.08 59.41 55.14 51.29 55.08 52.92 59.41 50.77 51.05
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 60.75 56.00 61.39 54.97 52.02 56.92 52.10 60.04 51.49 51.69
+Score-level Average 65.42 56.31 49.50 52.90 50.49 52.00 51.08 53.47 50.24 50.18
+Concat 54.21 52.62 58.42 51.60 50.81 54.15 50.77 55.45 50.47 50.52
+SE-Concat 65.42 56.31 66.34 52.03 51.05 54.46 50.96 56.64 50.60 50.81
+Cross-Atten 63.55 56.62 66.34 52.38 51.29 54.46 50.50 56.84 51.20 51.08
+MLP-Mixer 64.49 57.85 62.38 52.20 51.05 56.00 50.46 57.43 51.08 51.38
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 67.29 57.85 64.36 55.83 52.35 56.31 51.69 58.42 53.27 52.30
+Score-level Average 49.53 56.00 63.37 52.03 51.38 53.38 50.35 53.26 51.40 50.89
+Concat 57.01 50.15 58.42 52.29 50.97 56.25 50.15 53.47 52.34 51.05
+SE-Concat 57.01 49.85 63.37 52.38 50.57 56.62 50.77 54.46 53.64 51.29
+Cross-Atten 63.55 51.69 64.36 53.07 51.05 56.31 51.69 55.35 52.92 51.78
+MLP-Mixer 59.81 55.08 69.31 54.36 56.85 53.12 51.38 55.42 52.62 52.67
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 69.16 58.15 64.36 55.40 51.46 57.23 53.85 61.39 54.21 53.32
+Score-level Average 57.01 53.54 61.39 50.82 51.38 52.92 52.31 58.22 51.60 52.04
+Concat 56.07 54.15 60.40 51.69 50.16 53.23 52.00 58.32 51.40 52.00
+SE-Concat 59.81 54.77 52.48 52.03 51.78 56.95 53.65 59.45 51.50 52.40
+Cross-Atten 60.75 56.31 60.40 53.24 51.94 57.33 54.15 59.41 52.30 52.34
+MLP-Mixer 64.49 56.00 60.40 54.11 51.54 57.54 54.46 60.42 52.31 53.27
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 67.29 56.00 62.38 55.32 52.91 58.77 55.08 63.37 53.85 53.54
+Modality & Inputs Fusion Postion Fusion Method E to R E to B1 E to B2 E to M E to D Avg
+Score-level Average 49.68 51.21 56.46 49.53 53.50 53.05
+Concat 50.73 51.49 56.33 50.10 52.38 52.84
+SE-Concat 50.16 52.48 57.43 50.25 53.24 53.71
+Cross-Atten 50.65 53.46 57.83 50.47 53.07 54.22
+MLP-Mixer 51.38 55.42 57.25 51.24 54.11 54.35
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 52.27 56.21 60.09 51.45 54.54 55.78
+Score-level Average 49.84 52.46 57.87 50.03 52.72 52.83
+Concat 51.21 52.69 58.72 50.33 51.34 52.71
+SE-Concat 50.97 53.62 59.41 50.54 52.81 53.65
+Cross-Atten 50.57 54.99 59.70 50.67 52.72 54.25
+MLP-Mixer 51.21 55.08 60.08 50.95 52.20 54.80
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.70 55.45 61.39 51.63 54.19 55.79
+Score-level Average 50.24 54.50 57.43 50.65 51.08 52.74
+Concat 50.32 54.46 58.42 51.70 51.86 52.74
+SE-Concat 50.81 54.66 58.22 51.38 51.96 53.16
+Cross-Atten 50.97 55.32 58.62 52.18 53.85 54.05
+MLP-Mixer 50.65 55.65 58.77 52.10 54.11 54.90
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.02 56.44 60.41 53.16 55.06 56.14
+Score-level Average 49.50 53.23 58.00 51.21 52.29 53.52
+Concat 50.50 53.47 58.92 51.13 52.90 53.96
+SE-Concat 50.47 54.27 59.11 51.29 54.02 54.13
+Cross-Atten 50.67 55.45 61.29 52.02 54.28 54.65
+MLP-Mixer 51.49 56.31 60.40 52.27 55.92 54.73
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.48 56.25 63.37 54.91 56.09 56.82
+TABLE VI: Fusion results for multi-to-single cross-domain accuracy (%) using domain-simultaneous sampling.
 Domain-Simultaneous
-Score-level Average 53.23 43.56 57.43 50.94 51.40 58.42 48.51 51.93
-Concat 52.00 56.44 60.40 51.88 51.40 62.38 57.43 55.99
-SE-Concat 55.69 58.42 44.55 52.50 56.07 58.42 58.42 54.87
-Cross-Atten 51.08 52.48 55.45 51.56 57.01 60.40 57.43 55.06
-MLP-Mixer 53.85 43.56 62.38 52.19 56.07 61.39 59.41 55.55
-Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
-Score-level Average 49.54 55.45 52.48 51.25 54.21 59.41 55.45 53.97
-Concat 49.54 54.46 49.50 53.75 44.86 63.37 54.46 52.85
-SE-Concat 52.00 58.42 42.57 51.56 54.21 58.42 63.37 54.36
-Cross-Atten 50.46 52.48 60.40 53.12 52.34 58.42 59.41 55.23
-MLP-Mixer 53.54 57.43 52.48 50.31 52.34 60.40 47.52 53.43
-Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
-Score-level Average 48.62 58.42 56.44 50.94 51.40 56.44 53.47 53.68
-Concat 48.31 53.47 56.44 50.31 56.07 63.37 57.43 55.06
-SE-Concat 49.85 50.50 47.52 52.19 49.53 57.43 61.39 52.63
-Cross-Atten 51.38 63.37 61.39 51.25 48.60 59.41 58.42 56.26
-MLP-Mixer 55.69 59.41 57.43 53.75 50.47 61.39 50.50 55.52
-Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
-Score-level Average 54.77 57.43 57.43 52.81 60.75 58.42 56.44 56.86
-Concat 54.46 54.46 47.52 53.12 53.27 61.39 56.44 54.38
-SE-Concat 53.23 55.45 54.46 52.50 59.81 56.44 58.42 55.76
-Cross-Atten 48.62 57.43 50.50 53.12 48.60 58.42 56.44 53.30
-MLP-Mixer 48.92 48.51 59.41 52.81 54.21 57.43 51.49 53.25
-Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
+Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&M to D R&M to E R&B1 to B2 R&B1 to M R&B1 to D R&B1 to E B1&M to R B1&M to B2 R&B1&M to B2
+Score-level Average 53.23 43.56 50.57 50.16 57.43 50.94 56.44 50.16 51.40 58.42 48.51
+Concat 52.00 56.44 51.29 50.86 60.40 51.88 57.04 51.70 51.40 62.38 57.43
+SE-Concat 55.69 58.42 52.08 52.84 44.55 52.50 56.87 51.13 56.07 58.42 58.42
+Cross-Atten 51.08 52.48 52.51 52.24 55.45 51.56 57.13 51.94 57.01 60.40 57.43
+MLP-Mixer 53.85 43.56 53.02 53.32 62.38 52.19 57.74 52.10 56.07 61.39 59.41
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.23 57.43 53.50 53.40 63.37 52.19 57.39 51.86 57.01 63.37 62.38
+Score-level Average 49.54 55.45 50.73 50.73 52.48 51.25 57.56 50.81 54.21 59.41 55.45
+Concat 49.54 54.46 50.97 50.92 49.50 53.75 56.96 51.05 44.86 63.37 54.46
+SE-Concat 52.00 58.42 51.54 52.00 42.57 51.56 57.22 51.46 54.21 58.42 63.37
+Cross-Atten 50.46 52.48 52.02 53.08 60.40 53.12 57.30 51.13 52.34 58.42 59.41
+MLP-Mixer 53.54 57.43 52.10 53.32 52.48 50.31 57.56 52.18 52.34 60.40 47.52
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 64.36 53.97 53.43 56.44 53.75 57.82 52.35 58.88 59.41 58.42
+Score-level Average 48.62 58.42 51.21 50.29 56.44 50.94 56.48 51.78 51.40 56.44 53.47
+Concat 48.31 53.47 51.86 50.44 56.44 50.31 57.13 51.29 56.07 63.37 57.43
+SE-Concat 49.85 50.50 53.87 52.32 47.52 52.19 57.42 52.02 49.53 57.43 61.39
+Cross-Atten 51.38 63.37 54.50 53.24 61.39 51.25 57.74 52.27 48.60 59.41 58.42
+MLP-Mixer 55.69 59.41 55.06 53.24 57.43 53.75 57.43 52.99 50.47 61.39 50.50
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.08 55.45 57.39 54.20 58.42 54.37 58.02 52.83 52.34 61.39 60.40
+Score-level Average 54.77 57.43 55.50 50.90 57.43 52.81 55.45 51.46 60.75 58.42 56.44
+Concat 54.46 54.46 55.52 50.82 47.52 53.12 55.85 51.78 53.27 61.39 56.44
+SE-Concat 53.23 55.45 56.00 51.60 54.46 52.50 56.33 51.05 59.81 56.44 58.42
+Cross-Atten 48.62 57.43 56.26 52.18 50.50 53.12 57.46 51.62 48.60 58.42 56.44
+MLP-Mixer 48.92 48.51 56.61 53.68 59.41 52.81 58.40 52.21 54.21 57.43 51.49
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 52.31 59.41 57.30 54.57 58.42 52.50 59.41 52.84 58.88 61.39 60.40
+Modality & Inputs Fusion Postion Fusion Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Score-level Average 50.66 53.56 52.50 52.34 49.96 50.47 52.29 50.44 53.24 51.19 51.18
+Concat 51.38 54.46 50.31 52.31 53.33 52.18 53.40 51.25 53.24 51.19 51.18
+SE-Concat 51.49 54.86 51.85 52.20 55.49 53.07 55.15 51.88 54.87 54.81 53.94
+Cross-Atten 52.00 55.45 52.06 53.20 56.70 54.96 55.45 52.31 55.31 56.56 54.44
+MLP-Mixer 52.48 56.42 52.94 53.27 56.96 56.44 56.44 53.13 57.24 57.98 55.16
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.47 59.01 53.62 54.21 58.43 57.10 57.54 53.75 60.40 60.99 56.84
+Score-level Average 50.50 53.47 51.25 52.64 54.80 50.21 52.47 50.62 53.94 55.00 52.98
+Concat 52.34 53.87 50.62 52.98 56.78 53.48 54.26 51.55 53.69 56.16 53.21
+SE-Concat 52.10 55.15 51.38 53.85 58.08 55.75 54.35 53.25 54.06 56.26 54.14
+Cross-Atten 53.84 56.44 52.94 54.61 58.17 57.39 55.20 53.69 55.13 57.15 54.99
+MLP-Mixer 53.18 56.74 53.38 54.35 58.30 58.42 57.37 54.06 59.94 58.22 54.91
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.24 59.41 53.50 55.14 58.53 58.65 58.92 54.88 62.38 59.35 57.12
+Score-level Average 52.16 56.44 51.81 53.50 52.72 50.14 57.00 50.94 53.62 54.98 53.28
+Concat 52.08 56.04 51.20 53.44 56.87 50.77 57.50 51.57 54.69 57.39 54.18
+SE-Concat 53.21 57.27 52.75 54.36 57.13 56.09 58.51 52.69 55.75 58.22 54.29
+Cross-Atten 53.85 59.64 52.38 54.45 57.74 56.44 58.49 53.38 57.25 59.61 55.93
+MLP-Mixer 54.14 60.40 53.56 55.64 58.25 57.03 59.41 54.50 59.38 59.04 56.13
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.11 62.38 54.02 56.28 58.48 57.65 60.19 54.64 60.50 60.13 57.11
+Score-level Average 53.85 54.46 52.69 53.21 56.27 50.84 54.13 50.93 53.13 57.04 54.66
+Concat 55.42 57.43 52.62 54.97 56.61 53.33 54.87 51.50 54.06 57.54 54.43
+SE-Concat 56.44 58.13 52.90 56.07 56.88 56.18 55.02 53.16 54.44 57.22 55.32
+Cross-Atten 57.43 59.41 53.20 56.21 58.22 57.22 57.98 53.29 56.93 58.13 55.17
+MLP-Mixer 56.46 60.34 54.00 57.07 58.42 58.08 59.63 54.31 59.75 59.44 55.77
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.48 63.17 54.35 58.88 59.15 58.96 60.42 54.56 62.58 60.42 57.97
+TABLE VII: Fusion results for multi-to-single cross-domain accuracy (%) using domain-alternating sampling.
 Domain-Alternating
-Score-level Average 55.38 57.43 58.42 50.94 57.94 50.50 61.39 56.00
-Concat 55.08 58.42 63.37 51.56 62.62 62.38 53.47 58.13
-SE-Concat 50.46 49.50 63.37 52.19 57.01 65.35 56.44 56.33
-Cross-Atten 55.08 59.41 60.40 52.50 48.60 58.42 61.39 56.54
-MLP-Mixer 57.54 63.37 63.37 50.94 65.42 52.48 56.44 58.51
-Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.62 61.39 63.37 51.56 58.88 62.38 58.42 58.95
-Score-level Average 50.15 58.42 60.40 50.62 51.40 60.40 58.42 55.69
-Concat 53.23 62.38 54.46 51.25 59.81 54.46 54.46 55.72
-SE-Concat 50.15 52.48 64.36 50.31 59.81 55.45 50.50 54.72
-Cross-Atten 52.92 57.43 49.50 52.81 65.42 53.47 66.34 56.84
-MLP-Mixer 51.69 58.42 57.43 51.25 62.62 58.42 56.44 56.61
-Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.23 63.37 56.44 51.25 63.55 59.41 56.44 57.67
-Score-level Average 51.38 56.44 54.46 50.00 64.49 61.39 49.50 55.38
-Concat 49.85 62.38 58.42 51.88 61.68 55.45 51.49 55.88
-SE-Concat 50.46 62.38 58.42 50.00 57.94 59.41 65.35 57.71
-Cross-Atten 49.85 62.38 50.50 53.44 58.88 56.44 56.44 55.42
-MLP-Mixer 50.15 55.45 58.42 50.31 58.88 61.39 57.43 56.00
-Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 50.15 59.41 50.50 52.19 58.88 60.40 61.39 56.13
-Score-level Average 56.00 52.48 59.41 51.88 51.40 60.40 56.44 55.43
-Concat 51.69 58.42 58.42 52.81 57.01 61.39 56.44 56.60
-SE-Concat 56.31 59.41 58.42 50.31 53.27 61.39 55.45 56.37
-Cross-Atten 49.54 67.33 54.46 51.56 66.36 52.48 60.40 57.45
-MLP-Mixer 50.15 63.37 60.40 51.56 62.62 61.39 61.39 58.70
-Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 66.34 60.40 51.88 59.81 61.39 62.38 59.13
+Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&M to D R&M to E R&B1 to B2 R&B1 to M R&B1 to D R&B1 to E B1&M to R B1&M to B2 R&B1&M to B2
+Score-level Average 55.38 57.43 50.67 50.32 58.42 50.94 55.32 50.22 57.94 50.50 61.39
+Concat 55.08 58.42 50.29 50.54 63.37 51.56 56.98 52.73 62.62 62.38 53.47
+SE-Concat 50.46 49.50 51.18 52.78 63.37 52.19 56.80 52.05 57.01 65.35 56.44
+Cross-Atten 55.08 59.41 52.41 53.01 60.40 52.50 57.71 52.13 48.60 58.42 61.39
+MLP-Mixer 57.54 63.37 53.23 53.22 63.37 50.94 57.82 52.24 65.42 52.48 56.44
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 56.62 61.39 53.60 53.50 63.37 51.56 57.30 52.65 58.88 62.38 58.42
+Score-level Average 50.15 58.42 50.65 50.87 60.40 50.62 57.88 50.06 51.40 60.40 58.42
+Concat 53.23 62.38 51.11 50.95 54.46 51.25 57.04 51.24 59.81 54.46 54.46
+SE-Concat 50.15 52.48 51.66 52.55 64.36 50.31 57.73 51.26 59.81 55.45 50.50
+Cross-Atten 52.92 57.43 52.13 53.89 49.50 52.81 57.82 52.08 65.42 53.47 66.34
+MLP-Mixer 51.69 58.42 52.15 53.99 57.43 51.25 57.68 52.11 62.62 58.42 56.44
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.23 63.37 54.01 53.53 56.44 51.25 57.85 52.32 63.55 59.41 56.44
+Score-level Average 51.38 56.44 51.20 50.38 54.46 50.00 56.05 51.74 64.49 61.39 49.50
+Concat 49.85 62.38 51.00 50.48 58.42 51.88 57.23 51.86 61.68 55.45 51.49
+SE-Concat 50.46 62.38 53.24 53.84 58.42 50.00 57.72 52.34 57.94 59.41 65.35
+Cross-Atten 49.85 62.38 54.51 53.92 50.50 53.44 57.81 52.88 58.88 56.44 56.44
+MLP-Mixer 50.15 55.45 54.59 53.49 58.42 50.31 58.03 53.01 58.88 61.39 57.43
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 50.15 59.41 54.99 54.31 50.50 52.19 58.25 53.06 58.88 60.40 61.39
+Score-level Average 56.00 52.48 55.22 50.92 59.41 51.88 55.06 50.16 51.40 60.40 56.44
+Concat 51.69 58.42 55.38 51.09 58.42 52.81 55.32 51.25 57.01 61.39 56.44
+SE-Concat 56.31 59.41 56.22 51.28 58.42 50.31 56.38 51.99 53.27 61.39 55.45
+Cross-Atten 49.54 67.33 56.43 52.39 54.46 51.56 57.06 52.89 66.36 52.48 60.40
+MLP-Mixer 50.15 63.37 57.00 53.33 60.40 51.56 58.03 53.09 62.62 61.39 61.39
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 66.34 57.38 55.02 60.40 51.88 59.82 53.86 59.81 61.39 62.38
+Modality & Inputs Fusion Postion Fusion Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Score-level Average 50.65 53.45 52.35 51.16 48.96 50.67 52.31 50.35 54.34 52.34 53.10
+Concat 51.42 54.44 52.33 51.34 54.32 52.31 53.31 51.29 54.11 54.39 54.60
+SE-Concat 51.48 54.38 52.06 52.45 56.03 53.28 54.29 51.85 54.65 54.82 54.40
+Cross-Atten 52.34 55.26 53.10 53.52 56.13 54.65 54.99 53.05 55.33 55.69 55.01
+MLP-Mixer 52.46 55.39 53.15 53.65 56.35 56.33 56.05 53.38 56.89 57.82 56.07
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.89 57.02 53.45 54.62 58.16 57.01 58.24 54.02 59.89 59.85 56.94
+Score-level Average 50.53 52.88 51.38 52.63 54.28 51.04 52.85 51.25 53.22 55.68 53.57
+Concat 51.45 53.66 52.36 52.89 56.35 53.28 54.39 52.22 53.31 56.83 54.15
+SE-Concat 52.11 55.02 52.39 53.08 58.04 54.99 54.78 53.48 54.09 56.89 54.34
+Cross-Atten 52.98 56.36 53.24 53.96 58.26 56.84 55.76 53.87 55.68 57.63 55.62
+MLP-Mixer 53.42 56.79 53.33 54.21 58.45 57.05 57.35 54.33 59.88 58.47 55.98
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.64 58.98 53.06 54.76 58.95 58.28 58.92 55.21 62.05 59.86 56.96
+Score-level Average 52.17 56.42 51.28 52.36 53.15 50.58 57.08 51.43 53.34 54.27 53.79
+Concat 52.89 56.63 51.64 53.34 56.24 50.87 57.06 52.11 54.99 57.25 54.50
+SE-Concat 53.42 57.06 52.03 54.68 57.03 56.92 58.45 52.98 55.56 58.62 56.07
+Cross-Atten 53.22 58.99 52.65 54.92 57.05 56.83 58.69 53.89 57.84 59.37 55.74
+MLP-Mixer 54.65 59.78 53.08 55.01 58.34 57.36 59.05 54.54 59.08 59.38 56.26
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.38 61.49 54.56 56.38 58.65 57.82 60.18 55.38 61.05 61.20 56.89
+Score-level Average 52.85 56.84 52.41 53.04 56.48 51.22 54.28 51.02 53.45 57.39 54.21
+Concat 54.86 56.98 52.38 54.02 56.87 53.23 54.62 52.24 54.89 57.26 55.07
+SE-Concat 55.53 57.32 53.06 55.93 56.92 56.35 55.23 53.26 56.53 58.02 55.65
+Cross-Atten 56.88 59.63 53.66 55.89 57.98 57.26 57.01 53.82 56.87 58.25 56.58
+MLP-Mixer 56.08 61.02 55.24 57.46 58.35 58.35 59.55 54.38 57.05 59.36 57.58
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.23 63.93 55.05 58.95 59.65 59.01 61.14 55.01 59.96 61.35 58.63
+TABLE VIII: Fusion results for multi-to-single cross-domain accuracy (%) using domain-by-domain sampling.
 Domain-by-Domain
-Score-level Average 54.77 46.53 59.41 50.62 57.94 59.41 58.42 55.30
-Concat 53.85 53.47 63.37 51.88 57.94 65.35 58.42 57.75
-SE-Concat 56.62 44.55 63.37 52.19 54.21 58.42 59.41 55.54
-Cross-Atten 54.15 49.50 62.38 54.69 57.01 60.40 63.37 57.36
-MLP-Mixer 57.54 62.38 71.29 51.25 54.21 62.38 62.38 60.20
-Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 57.54 58.42 63.37 52.19 69.16 60.40 64.36 60.78
-Score-level Average 52.00 50.50 58.42 51.56 57.94 61.39 60.40 56.03
-Concat 50.46 59.41 62.38 52.19 60.75 56.44 57.43 57.01
-SE-Concat 55.38 48.51 58.42 52.81 62.62 57.43 59.41 56.37
-Cross-Atten 52.31 53.47 58.42 51.25 64.49 58.42 57.43 56.54
-MLP-Mixer 53.23 60.40 62.38 52.19 61.68 58.42 58.42 58.10
-Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.85 66.34 63.37 53.75 58.88 64.36 63.37 60.56
-Score-level Average 48.62 59.41 51.49 51.25 60.75 62.38 63.37 56.75
-Concat 57.23 60.40 56.44 51.25 56.07 57.43 56.44 56.47
-SE-Concat 55.08 60.40 59.41 54.06 60.75 61.39 57.43 58.36
-Cross-Atten 50.15 56.44 61.49 57.50 59.81 64.36 58.42 58.31
-MLP-Mixer 52.62 56.44 58.42 52.50 60.75 61.39 58.42 57.22
-Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 56.44 57.43 54.06 66.36 64.36 60.40 58.68
-Score-level Average 53.85 51.49 65.35 53.75 58.88 57.43 58.42 57.02
-Concat 52.92 52.48 55.45 53.75 62.62 62.38 56.44 56.58
-SE-Concat 51.38 60.40 58.42 51.56 57.01 62.38 59.41 57.22
-Cross-Atten 51.80 52.48 62.38 53.12 59.81 64.36 60.40 57.76
-MLP-Mixer 53.85 55.45 60.40 50.94 59.81 64.36 60.40 57.89
-Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 61.39 59.41 55.00 58.88 59.41 60.40 58.60
+Modality & Inputs Fusion Postion Fusion Method R&M to B1 R&M to B2 R&M to D R&M to E R&B1 to B2 R&B1 to M R&B1 to D R&B1 to E B1&M to R B1&M to B2 R&B1&M to B2
+Score-level Average 54.77 46.53 50.73 50.45 59.41 50.62 55.68 50.43 57.94 59.41 58.42
+Concat 53.85 53.47 50.37 51.02 63.37 51.88 56.32 52.39 57.94 65.35 58.42
+SE-Concat 56.62 44.55 51.23 52.39 63.37 52.19 56.56 53.56 54.21 58.42 59.41
+Cross-Atten 54.15 49.50 52.46 53.34 62.38 54.69 56.98 53.21 57.01 60.40 63.37
+MLP-Mixer 57.54 62.38 53.53 53.45 71.29 51.25 57.59 53.49 54.21 62.38 62.38
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 57.54 58.42 53.62 54.32 63.37 52.19 58.04 53.78 69.16 60.40 64.36
+Score-level Average 52.00 50.50 50.32 50.87 58.42 51.56 56.98 50.42 57.94 61.39 60.40
+Concat 50.46 59.41 51.24 50.95 62.38 52.19 57.25 51.28 60.75 56.44 57.43
+SE-Concat 55.38 48.51 51.35 52.54 58.42 52.81 57.83 51.87 62.62 57.43 59.41
+Cross-Atten 52.31 53.47 52.37 54.68 58.42 51.25 57.39 52.33 64.49 58.42 57.43
+MLP-Mixer 53.23 60.40 52.45 54.08 62.38 52.19 57.89 52.39 61.68 58.42 58.42
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.85 66.34 54.32 54.89 63.37 53.75 57.96 53.04 58.88 64.36 63.37
+Score-level Average 48.62 59.41 51.28 50.88 51.49 51.25 56.83 51.28 60.75 62.38 63.37
+Concat 57.23 60.40 51.35 50.39 56.44 51.25 57.38 51.38 56.07 57.43 56.44
+SE-Concat 55.08 60.40 53.86 52.83 59.41 54.06 57.92 52.98 60.75 61.39 57.43
+Cross-Atten 50.15 56.44 54.89 53.58 61.49 57.50 57.85 52.96 59.81 64.36 58.42
+MLP-Mixer 52.62 56.44 54.78 53.86 58.42 52.50 58.96 53.87 60.75 61.39 58.42
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 56.44 55.06 54.87 57.43 54.06 58.91 53.48 66.36 64.36 60.40
+Score-level Average 53.85 51.49 54.76 50.38 65.35 53.75 55.64 50.32 58.88 57.43 58.42
+Concat 52.92 52.48 55.93 51.77 55.45 53.75 55.84 51.28 62.62 62.38 56.44
+SE-Concat 51.38 60.40 56.87 52.36 58.42 51.56 56.92 51.98 57.01 62.38 59.41
+Cross-Atten 51.80 52.48 56.82 53.58 62.38 53.12 57.83 52.38 59.81 64.36 60.40
+MLP-Mixer 53.85 55.45 57.82 53.69 60.40 50.94 58.92 53.56 59.81 64.36 60.40
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 61.39 58.05 55.72 59.41 55.00 60.31 54.34 58.88 59.41 60.40
+Modality & Inputs Fusion Postion Fusion Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Score-level Average 50.76 53.62 52.31 51.22 49.62 50.38 52.35 51.15 54.37 52.12 52.98
+Concat 51.25 54.21 52.38 51.42 53.21 52.46 53.46 52.31 54.88 54.38 54.49
+SE-Concat 51.36 54.38 52.36 52.36 54.87 53.49 54.63 52.33 55.62 55.72 54.27
+Cross-Atten 52.45 55.32 53.67 53.34 55.74 54.42 56.67 53.87 55.71 55.84 55.45
+MLP-Mixer 52.66 55.48 53.48 53.76 56.72 56.73 56.72 54.83 56.82 55.71 56.78
+Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.86 57.34 53.86 54.11 58.24 57.02 57.39 54.87 58.78 58.72 57.59
+Score-level Average 50.38 53.25 52.33 52.39 54.42 51.25 52.38 51.24 54.31 53.43 53.63
+Concat 51.49 53.48 52.47 53.18 56.31 53.32 54.32 52.33 54.33 56.72 54.68
+SE-Concat 52.95 55.28 52.38 53.91 57.82 54.75 55.64 54.32 54.21 56.83 55.06
+Cross-Atten 52.86 56.83 53.34 53.78 58.62 57.33 55.87 54.65 55.53 56.79 55.61
+MLP-Mixer 53.47 56.74 53.48 54.27 58.72 57.38 57.64 54.87 55.49 57.83 56.35
+Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.87 58.69 55.87 54.86 58.98 58.24 58.99 55.68 61.39 58.99 58.13
+Score-level Average 52.86 56.31 51.85 52.38 53.21 51.27 57.67 51.23 54.06 54.72 54.43
+Concat 52.65 55.89 51.64 53.34 56.72 52.23 57.83 52.35 54.65 55.49 54.69
+SE-Concat 53.49 57.28 52.38 54.96 57.46 56.93 58.72 52.39 55.68 57.89 56.35
+Cross-Atten 53.64 58.72 53.03 54.73 57.68 56.73 58.31 53.48 57.83 59.83 56.73
+MLP-Mixer 54.67 59.81 53.48 55.92 58.92 57.84 59.98 54.63 58.75 59.86 56.95
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.78 60.45 54.72 56.03 59.21 58.82 60.12 55.69 60.33 62.31 57.89
+Score-level Average 52.38 57.33 52.88 53.42 56.33 52.31 54.32 51.25 53.46 58.93 54.90
+Concat 53.08 58.31 53.02 54.33 56.31 53.42 54.76 52.38 54.55 58.72 55.23
+SE-Concat 54.89 58.39 53.76 55.41 57.35 56.65 55.65 54.83 56.73 58.79 56.24
+Cross-Atten 55.98 59.74 53.47 55.64 57.36 57.82 57.86 55.73 56.79 58.22 56.84
+MLP-Mixer 56.73 60.28 55.08 58.02 58.72 58.93 59.93 55.76 58.32 59.89 57.66
+Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.47 63.48 56.71 58.36 59.55 60.33 61.35 57.83 60.36 62.35 58.88
 4.5 Cross-domain Testing with Multimodal Fusion
@@
 The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram).
-It results in four combinations of either two or three types of inputs.
-As shown in Table IV, among all the fusion methods, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results on average.
+As shown in Table V, there are 25 domain transfer tasks, and the proposed Atten-Mixer consistently outperforms other methods across all modalities and fusion types, achieving the highest average accuracies.
+This demonstrates its strong ability to learn transferable representations.
+Feature-level fusion generally yields better performance than score-level fusion, highlighting the benefit of early interaction between modalities.
+Among the modalities, the best average performance across all the datasets is achieved when fused with Atten-Mixer.
 Multi-to-Single Domain with Fusion.
-We benchmarked multi-to-single cross-domain generalization with different fusion methods by three types of cross-testing strategies.
-In total, seven sub-experiments were conducted for different domain combinations.
-The results are shown in Tabel V.
-On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively.
-Among these, for both domain-simultaneous and domain-by-domain strategies, the best results were achieved by taking visual (face frames) and visual (AU+gaze+affect) features as input, while the best result for domain-alternating was achieved by using visual (face frames), visual (AU+gaze+affect) and audio (Mel spectrogram) features.
-Taking a close look at the average fusion results, for each type of modality input, Atten-Mixer achieved the best among the six fusion methods.
-showing the effectiveness of the proposed method.
-By using Atten-Mixer, in general, the average result showed slightly better when using the domain-by-domain strategy and taking visual (face frames) and visual (AU+gaze+affect) features as input.
-To sum up, the results showed that on the current publicly available datasets, visual features were better when it came to multi-to-single cross-domain generalizability.
-However, the performance differences were not a large gap, and there were also no significant differences by comparing two-to-one domain and three-to-one domain cross-testing performances.
+Tables VIVIII present multi-to-single domain generalization results across 21 transfer tasks using four different modality combinations under three sampling strategies.
+Across all settings, the proposed Atten-Mixer consistently outperforms baseline fusion methods, achieving the highest average accuracies.
+Notably, feature-level fusion of Visual (AU+Gaze+Affect) + Audio using Atten-Mixer reaches the best overall average performance: 57.97% (Simultaneous), 56.35% (Alternating), and 58.88% (domain-by-domain), showing its strong cross-domain generalization.
+Among different strategies, domain-by-domain consistently yields the highest averages across modalities, similar to the observation in the multi-to-single domain generalization results without fusion methods.
+Besides, feature-level fusion generally outperforms score-level fusion, indicating the importance of deep modality interaction for fusion.
+Additionally, multimodal combinations (e.g., Visual + Audio) consistently outperform unimodal inputs, validating the effectiveness of multimodal fusion methods on visual and audio cues in deception detection.
+Overall, the results show the effectiveness of Atten-Mixer and underscore the advantage of modality fusion in audio-visual deception detection.
+TABLE IX: The fusion results of multi-to-single cross-domain generalization accuracy (%) for multimodal inter-domain gradient matching (MM-IDGM).
+The numbers in the brackets show the improved accuracies compared with Atten-Mixer.
+Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
+Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 53.50 53.40 57.39 51.86
+IDGM 55.22 57.45 59.85 51.41 61.34 59.41 60.61 52.70 52.99 57.05 51.22
+Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 57.85 58.42 60.40 52.50 65.22 62.05 61.42 53.88 54.02 58.51 52.87
+Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 53.97 53.43 57.82 52.35
+IDGM 54.43 60.40 61.07 51.40 59.21 59.88 59.75 53.02 53.11 56.24 52.36
+Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.54 59.41 62.38 51.56 59.07 62.38 59.50 54.26 54.29 58.33 53.46
+Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 57.39 54.20 58.02 52.83
+IDGM 55.49 54.92 58.36 50.31 53.15 60.35 59.21 56.87 53.40 57.54 52.33
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 56.05 56.03 59.37 52.41 53.40 62.42 60.98 58.24 55.06 58.72 51.24
+Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.30 54.57 59.41 52.84
+IDGM 52.45 57.02 60.42 54.25 58.42 61.79 60.03 58.02 54.21 59.45 53.14
+Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 52.77 58.43 61.39 55.57 59.46 62.58 60.60 58.75 55.87 60.85 53.01
+Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
+Atten-Mixer(Ours) 53.47 59.01 53.62 54.21 58.43 57.10 57.54 53.75 60.40 60.99 56.84
+IDGM 54.12 58.04 54.36 58.92 59.03 57.84 57.93 54.46 60.24 60.54 56.89
+Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 54.88 59.26 54.38 59.34 58.28 58.28 58.66 55.25 60.34 60.58 57.92 (+1.08)
+Atten-Mixer(Ours) 54.24 59.41 53.50 55.14 58.53 58.65 58.92 54.88 62.38 59.35 57.12
+IDGM 53.29 59.88 54.01 55.39 58.98 59.35 58.43 55.34 62.57 60.42 57.07
+Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.56 59.81 55.36 56.03 59.21 59.32 59.42 55.23 61.92 62.35 57.64 (+0.52)
+Atten-Mixer(Ours) 55.11 62.38 54.02 56.28 58.48 57.65 60.19 54.64 60.50 60.13 57.11
+IDGM 55.87 61.24 55.02 56.21 57.03 57.95 60.25 55.32 60.61 59.85 56.73
+Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 55.99 62.68 55.36 57.02 59.31 58.34 60.58 55.41 60.89 61.78 57.68 (+0.57)
+Atten-Mixer(Ours) 57.48 63.17 54.35 58.88 59.15 58.96 60.42 54.56 62.58 60.42 57.97
+IDGM 56.98 62.11 54.68 58.25 59.64 59.31 60.15 54.73 61.99 61.23 58.01
+Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 58.13 64.22 55.05 59.31 60.14 59.46 61.88 56.93 62.64 62.34 59.02 (+1.05)
+Performance of MM-IDGM We proposed MM-IDGM to enhance multi-to-single domain generalization in deception detection.
+Table IX presents the experimental results under the domain-simultaneous sampling strategy, comparing Atten-Mixer (baseline), IDGM, and the proposed MM-IDGM across various modality combinations.
+In all tested configurations, MM-IDGM consistently outperforms Atten-Mixer, achieving up to +1.08% improvement in average accuracy, particularly in the case of Visual (Face Frames) + Visual (AU+Gaze+Affect).
+This demonstrates the effectiveness of aligning gradient directions across domains at the feature level in a multimodal setting.
+While IDGM improves performance in many cases, it occasionally degrades accuracy.
+This may be due to conflicts from aligning gradients independently per modality.
+In contrast, MM-IDGM aligns all modalities jointly, yielding more stable and transferable features, and showing promising improvements in multimodal cross-domain generalization.
+However, its effectiveness can vary depending on the complexity of the target domain, data quality, and the severity of distribution shifts.
 Figure 4: Ablation study for attention-mixer layers.
 The number of layers 4, 5, 6, and 7 are compared.
-The modality and inputs for A, B, C, and D are in line with those in Table IV from the top to the bottom.
+The modality and inputs for A, B, C, and D are in line with those in Table V from the top to the bottom.
 Ablation Study for Attention-Mixer Fusion Module.
@@
 4, the number of attention-mixer layers was set to 4, 5, 6, and 7.
-The modalities and inputs in Table IV are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram).
+The modalities and inputs in Table V are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram).
 The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
-TABLE VI: The fusion results of multi-to-single cross-domain generalization accuracy (%) for multimodal inter-domain gradient matching (MM-IDGM).
-The numbers in the brackets show the improved accuracies compared with Atten-Mixer.
-Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
-Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
-IDGM 55.22 57.45 59.85 51.41 61.34 59.41 60.61 57.90
-Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 57.85 58.42 60.40 52.50 65.22 62.05 61.42 59.69 (+1.26)
-Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
-IDGM 54.43 60.40 61.07 51.40 59.21 59.88 59.75 58.02
-Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.54 59.41 62.38 51.56 59.07 62.38 59.50 58.26 (+0.12)
-Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
-IDGM 55.49 54.92 58.36 50.31 53.15 60.35 59.21 55.97
-Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 56.05 56.03 59.37 52.41 53.40 62.42 60.98 57.23 (+0.45)
-Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
-IDGM 52.45 57.02 60.42 54.25 58.42 61.79 60.03 57.77
-Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 52.77 57.43 61.39 55.57 59.46 62.58 60.60 58.54 (+0.92)
-Performance of MM-IDGM We proposed MM-IDGM and conducted experiments for multi-to-single cross-domain generalization, with the results presented in Table VI.
-Both the IDGM and MM-IDGM algorithms are developed based on the domain simultaneous sampling method.
-Using the proposed Atten-Mixer fusion method as the baseline, we applied both the original IDGM and the proposed MM-IDGM.
-The results show that MM-IDGM consistently improves average accuracies across different modalities and inputs.
-However, cross-domain generalization in deception detection remains highly challenging, as some accuracies even decreased after applying IDGM.
-This may be due to the difficulty in optimizing the inner product between different domains.
-While MM-IDGM may not always perform optimally in certain cases, such as testing on the B2 dataset where dataset quality affects the results, it generally shows better performance in cross-domain generalization.
+The results showed that deeper fusion networks (such as 6 layers vs.
+4 or 5 layers) improve generalization by enhancing cross-modal interactions, but overly deep architectures (e.g., 7 layers) may bring diminishing returns.
+Therefore, it is important to carefully select the fusion depth to balance representation capacity and the risk of overfitting.
 4.6 Network Details
-Here, we present the detailed network architectures in Table VII, including encoders, the fusion module, and classifiers.
+Here, we present the detailed network architectures in Table X, including encoders, the fusion module, and classifiers.
 This network architecture is utilized for experiments on multi-to-single cross-domain generalization with fusion.
-TABLE VII: Architecture details for different modalities and fusion module.
+TABLE X: Architecture details for different modalities and fusion module.
 Modality & Inputs: Visual (Face Frames)